### v3

# PAWS-VMK: A Unified Approach To Semi-Supervised Learning And Out-of-Distribution Detection

Abstract
This paper describes PAWS-VMK, a prototypical deep learning approach that obtains state-of-the-art results for image classification tasks in both a semi-supervised learning (SSL) and out-of-distribution (OOD) detection context. We consider developments in the fields of SSL, OOD detection, and computer vision foundation models to introduce a number of innovations that connect the key ideas within these works to create PAWS-VMK. These innovations include (1) parametric von Mises-Fisher Stochastic Neighbour Embedding (vMF-SNE) to initialise a projection head for SSL using the high-quality embeddings of the foundation model; (2) the PAWS-MixMatch loss, that creates more compact embeddings and obtains higher accuracy in comparison to the consistency loss used in PAWS and (3) simple $k$-Means prototype selection (SKMPS), a simple technique that obtains competitive performance with more complex unsupervised label selection approaches. PAWS-VMK sets new benchmarks in semi-supervised learning for CIFAR-10 (99.2%) and CIFAR-100 (89.8%) with four labelled instances per class, and Food-101 (90.1%) with two labelled instances per class. We also observe that PAWS-VMK can efficiently detect OOD samples in a manner that is competitive with specialised methods specifically designed for this purpose, achieving 93.1/98.0 and 95.2/96.3 on the CIFAR-10 and CIFAR-100 OpenOOD benchmarks.
1 Introduction
From managing weeds [1], to food safety [2], to diagnosing cancer [3], deep learning is used extensively in computer vision applications where specialised knowledge is needed to annotate data. Within these contexts, identifying out-of-distribution (OOD) data, such as plants or food contaminants not present in the training dataset, can also be of great importance. This creates a need for semi-supervised learning (SSL) approaches that can learn accurate models with only a small number of annotations, and that also perform well on detecting OOD examples. While SSL [4, 5] and OOD detection [6, 7] methods are well developed, few studies consider them simultaneously. While it is possible that separate models could be used for each task, different training approaches may lead to inconsistencies between them and using multiple models will increase deployment costs.
Prototype deep learning approaches, which learn a metric or semi-metric space for classifying items based on their proximity to class prototypes, present a promising solution for this challenge. Prototypical approaches have been previously used in semi-supervised learning [8] and OOD detection [9, 10]. However, these approaches use very different techniques to achieve their goals. For example, where the Predicting View-Assignments With Support Samples (PAWS) [8] SSL algorithm uses the labelled dataset to generate prototypes, the PrototypicAl Learning with a Mixture of prototypes (PALM) [10] OOD detection method uses an averaging approach to non-parametrically learn them during training. Notably, where PALM employs a specialised loss to ensure the in-distribution training data is compactly represented by well separated prototypes, PAWS uses a consistency based loss where labels are swapped between different views of an image to learn within an SSL context.
We also consider these methods in the context of recent developments in computer vision foundation models. These models are trained using large-scale self-supervised learning [11] or language-image pretraining [12] and map images into an expressive feature space where images with similar semantic information are grouped together, showing promising generalizability across multiple image domains [13, 11, 12]. Leveraging these foundation models, such as the DINOv2 family of ViT models [11], as backbone networks is a promising direction for both OOD detection and SSL approaches to improve performance as well as training efficiency.
However, incorporating the strengths of PAWS, PALM, and the DINOv2 foundation models into a single unified approach that performs better than the component parts on OOD detection and SSL is a non-trivial task. The consistency loss used within PAWS could lead to issues at decision boundaries, where different views are labelled as different classes and minimising the loss in this situation would lead to less sparse representations and poor performance on OOD detection tasks. Without any modifications, we have found that naively training PAWS with a frozen DINOv2 backbone can perform substantially worse than state-of-the-art results trained from scratch on small networks (Table 1). Conversely, naively incorporating elements of the PALM loss into PAWS adversely impacts SSL accuracy (Table S10).
To address these challenges, we propose PAWS-VMK a novel semi-supervised learning approach designed to be used with foundation models that sets new benchmarks in SSL and OOD detection. This method rectifies several challenges with the PAWS approach that we have identified to reduce accuracy and hurt OOD detection performance when a frozen foundation model is employed as the backbone network. Our main contributions are as follows:
  1. 1.
    We describe PAWS-VMK, a semi-supervised learning approach for image classification designed to be used with foundation models. This method is based on the PAWS algorithm, but incorporates the following innovations: (1) parametric vMF-SNE initialisation using the foundation model backbone to improve the quality of the projection head embeddings, (2) the PAWS-MixMatch loss which improves accuracy as well as improving OOD detection performance, and (3) the simple $k$-means prototype selection (SKMPS) strategy which outperforms more complex unsupervised label selection approaches within this context.
  2. 2.
    We demonstrate that PAWS-VMK obtains better accuracy compared to PAWS with the same foundation model backbone and hyper-parameters across a range of datasets. We also compare our results to the state-of-the-art and find the SSL performance of PAWS-VMK is competitive.
  3. 3.
    We demonstrate that PAWS-VMK outperforms PAWS on the CIFAR-10 and CIFAR-100 OpenOOD [7] benchmarks. We also compare our results to similar, supervised OOD detection approaches and find the performance of PAWS-VMK is competitive despite only using a small amount of labelled data.
2 Related work
Foundation models. In computer vision, foundation models can be used as functional maps from the image domain to a representation space that captures the semantic content of the image with minimal loss of information. The degree to which semantic content is retained can be tested by undertaking downstream tasks within the representation space, such as image classification, segmentation and object detection [14]. Current state-of-the-art foundation models include methods that combine several self-supervised approaches on large datasets, such as DINOv2 [11], and other approaches that use language-image pretraining approaches such as CLIP [12, 15].
Semi-supervised learning. In semi-supervised learning (SSL), a model is trained to perform a particular task, such as image classification, using a small set of labelled data and a much larger set of unlabelled data. By taking into account the distribution of the unlabelled data within the learning process, SSL methods can train more accurate models compared to approaches that only utilize the labelled data within this context [5]. The earliest SSL approaches used pseudolabelling [16], and a number of modern approaches have extended on this idea such as FixMatch [17], which employs image augmentation to efficiently propogate labels, and other similar methods [18, 19, 20].
Out-of-distribution detection. The problem of out-of-distribution (OOD) detection in deep learning was introduced when it was observed that neural networks tend to provide overconfident results when making predictions on images with previously unseen classes or that are unrecognizable [21]. OOD detection methods are designed to detect when inference is going to be made on such data, so an appropriate alternative action can be taken [22]. Various OOD detection approaches have been proposed [23, 24, 25, 7], including several methods such as PALM that use a prototypical approach to create embeddings in which the in-distribution images are tightly clustered [10, 9]. Most OOD detection methods assume class labels are available for all in-distribution images, and a few have considered the entirely unsupervised case [10].
Stochastic Neighbour Embedding. Given a high dimensional dataset, Stochastic Neighbour Embedding (SNE) [26] methods create a representation of this data in a lower dimensional space while attempting to ensure the relationship between items in the dataset is preserved. They are often used for plotting high dimensional data in two or three dimensional spaces [27]. One of the most well-known and commonly used methods is the t-distributed Stochastic Neighbour Embedding (t-SNE) approach [28].
Prototype selection strategies. The goal of prototype selection is to choose from an unlabelled dataset an optimal set of images with which to label and undertake semi-supervised, active or few-shot learning. Other works have referred to this challenge as unsupervised selective labelling [29], or the cold start problem [30]. The optimal image set is defined as the set of images for a given labelling budget that obtain maximal performance under a particular learning paradigm. While it has been found that using these prototypical image sets can lead to significant performance improvements [31, 17, 19], identifying them is challenging. The Unsupervised Selective Labelling (USL) [29] approach is the current state-of-the-art method for unsupervised prototype selection when using a frozen backbone, such as a foundation model or when using self-supervised learning.
3 Methodology
3.1 Preliminaries
Setup. Semi-supervised learning (SSL) aims to train a model using a small labelled dataset $D_{l}:=\{x_{l},y_{l}\}$ and a much larger unlabelled dataset $D_{u}:=\{x_{u}\}$. We simultaneously consider out of distribution (OOD) detection, where the goal is to train a model with in-distribution data $D_{l}\cup D_{u}$, and at test-time try to detect samples from OOD datasets. For this work, we only compare OOD samples to the small labelled set $D_{l}$ for OOD detection.
We are interested in training models to undertake these tasks using a frozen foundation model backbone. The foundation model maps the data onto a representational space, as defined by a neural network $f_{\theta}:X\in\mathbb{R}^{3\times H\times W}\to Z\in\mathbb{R}^{d_{1}}$. Here $H$ and $W$ are the height and width of the input images $X$, and $d_{1}$ is the dimensionality of this representation space. This representation space $Z$ can also be referred to as a latent space.
Predicting View Assignments With Support Samples. PAWS [8] is a prototypical SSL method that takes inspiration from self-supervised learning approaches. PAWS adapts the idea of prototypes from Swapping Assignments Between Multiple Views (SwAV) [32], which learns prototype vectors to cluster data for self-supervised learning. However, PAWS uses the embeddings of the labelled images $D_{l}$ as prototypes instead of learning them non-parametrically.
Architecture. PAWS [8] uses an architecture with a backbone $f_{\theta}$ and a projection head, defined by a second neural network $g_{\phi}:Z\in\mathbb{R}^{d_{1}}\to\hat{Z}\in\mathbb{R}^{d_{2}}$. This creates a second latent space $\hat{Z}$ with dimension $d_{2}$, which depending on the problem may be smaller or larger than $d_{1}$. Projection heads are often used in self-supervised learning [33].
Inference. PAWS [8] can make predictions using the projection head directly, by considering the cosine similarity between the normalized projection of the image into the latent space, $\hat{z}_{j}=\frac{g_{\phi}(f_{\theta}(x_{j}))}{||g_{\phi}(f_{\theta}(x_{j}))||}$, with the normalized projection of the labelled data $\hat{z}_{S}$
$p(y|x_{j},D_{l})$ $=\sigma\left(\hat{z}_{j}\cdot\hat{z}_{S}^{T}/\tau\right)y_{S}=p_{j}$ (1)
This can be derived by treating PAWS as a generative model [34] with a von Mises-Fischer kernel [35, 36], as described further in Appendix A within the supporting information.
Training objective. PAWS [8] uses a consistency loss as a supervision signal, employing cross-entropy to encourage the prediction of two views of the same image to be similar. The loss is given by
$\mathcal{L}_{\text{PAWS}}:=\frac{1}{2n}\sum_{i=1}^{n}\left(H(\rho(p_{i}^{+}),p_{i})+H(\rho(p_{i}),p_{i}^{+})\right)-H(\overline{p})$ (2)
where $p_{i}$ and $p_{i}^{+}$ are the probabilities given by Eq. 1 for different augmented views of an image, $n$ is the number of unlabelled images and $\rho$ is a sharpening function used to prevent representation collapse as will be discussed in Section 3.2.1. In the final term, $\overline{p}$ is the average of the sharpened predictions on the unlabelled images. This mean entropy maximization regularization (me-max) term is used to maximise the entropy of the average predictions [8]. This helps ensure that the association of the unlabelled points with each class is more uniform. An overview of the PAWS training framework is shown in Fig. 1(b).
Frozen backbone $f_{\theta}$ Augmented views of $x\in X$ $x^{+},x^{*}$ $z^{*},z^{+}\in Z$ $P$ Projection head $g_{\phi}$ $\hat{z}^{*},\hat{z}^{+}\in\hat{Z}$ vMF-SNE $L=KL(Q||P)$ SKMPS prototype selection // $Q$
(a) Initialisation framework. SKMPS uses a clustering approach on the backbone embeddings to select an optimal set of prototypes for training. vMF-SNE initialises the embeddings of the projection head to reflect the local structure of the backbone embeddings over the dataset.
Frozen backbone $f_{\theta}$ Augmented views of $x\in X$ $x^{+},x^{*}$ Projection head $g_{\phi}$ $\hat{z}^{*},\hat{z}^{+}\in\hat{Z}$ Labelled prototypes $x_{S},y_{S}$ Frozen backbone $f_{\theta}$ Projection head $g_{\phi}$ $\hat{z}_{S}\in\hat{Z},y_{S}$ $p^{*},p^{+}$ PAWS-MixMatch loss $H\left(\rho(\frac{p^{*}+p^{+}}{2}),p^{*}\right)$ Consistency loss $H\left(\rho(p^{+}),p^{*}\right)$ ////
(b) Training framework. PAWS-VMK uses the PAWS-MixMatch loss, while PAWS uses a consistency loss.
Figure 1: Overview of our proposed framework for PAWS-VMK.
3.2 Improving the PAWS approach
Our goal is to adapt the PAWS semi-supervised learning approach to perform well with a foundation model, and obtain better OOD detection results. To achieve this we propose PAWS-VMK, which uses the PAWS-MixMatch loss, a vMF-SNE initialisation strategy, and the SKMPS strategy for selecting optimal prototypes. An overview of these components is provided in Fig. 1.
3.2.1 Sharpening and the PAWS-MixMatch loss.
The consistency loss employed by PAWS in Eq. 2 penalises the prediction from one augmented view of an image differing from another. However, a sharpening function $\rho$ is used within this loss to prevent representational collapse and encourage confident predictions,
$\rho(p)_{i}$ $:=\frac{p_{i}^{1/T}}{\sum_{i=1}^{C}p_{i}^{1/T}}$ (3)
with the temperature parameter fixed at $T=0.25$ [8]. Fig. S6 in the supplementary material shows that for moderately confident predictions, this level of sharpening quickly leads to an entirely confident sharpened prediction, particularly with a large number of classes. If a particular image is augmented such that each view is sharpened to a confident prediction of a different class, this will cause undesirable mixing at the class boundaries. We observe for some datasets this results in performance degradation over a training run when this loss is used with frozen DINOv2 embeddings (Fig. S7).
Instead we propose to use a MixMatch [37] inspired loss, that averages the predictions from multiple views before sharpening to provide a more consistent supervision signal, improve accuracy and produce more compact embeddings to improve OOD detection results. The PAWS approach uses two global views within the consistency loss, and we propose to jointly sharpen the two global views together to create the PAWS-MixMatch loss
$\mathcal{L}_{\text{PAWS-MixMatch}}:=$ $\frac{1}{2n}\sum_{i=1}^{n}\left(H\left(\rho\left(\frac{p_{1}+p_{2}}{2}\right),p_{i}\right)\right.$
$+\left.H\left(\rho\left(\frac{p_{1}+p_{2}}{2}\right),p_{i}^{+}\right)\right)-H(\overline{p}).$ (4)
3.2.2 Initialising the projection head with vMF-SNE.
If we are using a pretrained backbone model $f_{\theta}$, which provides a good representation of our data in the first latent space $Z$, a good starting point for our projection head $g_{\phi}$ should reflect the local structure of the first latent space $Z$ within the second $\hat{Z}$. This is not a given if we randomly initialize the parameters of the projection head, and without placing restrictions on its architecture and the dimensionality of our latent spaces, we cannot simply initialise it as an identity mapping.
This problem is similar to knowledge distillation [38], but rather than aligning outputs onto a fixed set of classes, we are aligning outputs directly within an embedding space. We propose to instead use a Stochastic Neighbour Embedding (SNE) [26] approach. This method involves building two probability distributions $P,Q$ that describe the similarity of data points in the origin space $Z$ and the target space $\hat{Z}$. The idea is for points $i,j$ in $Z$, $p_{ij}\in P$ will have high probability if they are similar and low probability if they are dissimilar. We can then fit the target distribution $Q$ in the target space $\hat{Z}$ to the distribution in the origin space $P$ by minimizing the KL divergence
$\text{KL}(Q||P)=\sum_{ij}p_{ij}\log\frac{p_{ij}}{q_{ij}}$ (5)
where $p_{ij}$ and $q_{ij}$ describe the similarity between points in the origin and target space. Inspired by the relationship between inference within PAWS and von Mises-Fisher (vMF) kernels, we explore a novel parametric vMF-SNE approach as described in Appendix A within the supporting information to initialise the projection heads of our networks.
3.2.3 Prototype selection.
In this work we propose the simple $k$-means prototype selection (SKMPS) strategy. This approach provides for a diverse sampling of prototypes in three steps: (1) calculate normalized embeddings using a foundation model, then (2) cluster the data into $k$ clusters. We set $k$ to be the number of images (budget) desired for the labelled dataset $D_{l}$. And (3) select the image closest to the centroid of each cluster as a prototype.
We choose $k$-means clustering [39] for the clustering algorithm in the second step as it is fast to compute and allows for the selection of a fixed number of clusters. The $k$-means algorithm minimises the sum of squared euclidean distances to the cluster centroids [39]. By normalizing the image embeddings given by the foundation model in the first step, the squared euclidean distance becomes proportional to cosine distance, which is used for determining image similarity with foundation models [11, 12]. In this sense, the SKMPS procedure can be interpreted as identifying clusters of visually similar images, and selecting the most representative image from each cluster as a prototype.
4 Experiments
4.1 Implementation details
We closely follow the implementation of PAWS [8] for our experiments, but are able to shorten the training regime as we freeze the weights of the DINOv2 foundation models that are used for the model backbone. All of the semi-supervised learning models that are fitted in this paper were trained for 50 epochs, and the vMF-SNE initialisation runs were trained for 20 epochs. This meant that even for the Food-101 dataset with 75,000 images, training a single model took less than two hours with two Nvidia V100 32GB graphics cards for the ViT-S/14 and ViT-B/14 backbone, and around six hours for the ViT-L/14 backbone. In contrast to the original PAWS approach [8], we only use the projection head as a classifier, using the inference formula in Eq. 1 to make predictions using the labelled prototypes. Further details and hyperparameters are provided in the supplementary material.
4.2 Datasets
Seven datasets were selected to demonstrate our approach. These include CIFAR-10 (10 classes), CIFAR-100 (100 classes) [40], EuroSAT (10 classes) [41], Flowers-102 (102 classes) [42], Food-101 (101 classes) [43], Oxford Pets (37 classes) [44] and DeepWeeds (2 classes) [1]. While DeepWeeds has originally 9 classes, we treat it as a binary classification problem to identify weeds versus not weeds, due to the size of the negative class.
When sampling the prototypes for semi-supervised learning, we selected 4 (CIFAR-10), 4 (CIFAR-100), 4 (EuroSAT), 27 (DeepWeeds - binary), 2 (Flowers-102), 2 (Food-101) and 2 (Oxford Pets) images per class. Results are reported as the mean of five runs and the standard deviation, with the same sets of prototypes being used for all experiments unless specified otherwise, where the USL [29] or SKMPS strategy is used to select prototypes. For EuroSAT and DeepWeeds, where preset validation splits were not available, we sampled 500 images per class, and 30% of the images for the 9 original classes respectively.
For the OOD detection experiments, we follow the OpenOOD [7] benchmark in choosing comparable datasets that are similar (near OOD) to CIFAR-10 and CIFAR-100, which include each other respectively and the Tiny ImageNet [45] dataset. We also explore the dissimilar (far OOD) benchmark for these datasets, which include Places365[46], DTD [47], MNIST [48] and SVHN [49].
4.3 Main results
Table 1: Performance of PAWS-VMK versus PAWS and other methods. Comparable model sizes in each column are underlined, and the best performing model of these is bolded. Entries noted with * use a larger labelled dataset (1% of the data) than in our benchmarks.
Methods Backbone Datasets
C10 C100 Food
FreeMatch [50] WRN-28-2 1M 95.1
+SemiReward [18] ViT-S-P4-32 21M 84.4
Semi-ViT [20] ViT-Base 86M 82.1*
PAWS [8] DINOv2 ViT-S/14 (f) 21M 92.9±2.1 70.9±1.5 75.1±2.0
PAWS-VMK DINOv2 ViT-S/14 (f) 21M 95.9±0.0 77.3±0.4 81.7±0.7
DINOv2 ViT-B/14 (f) 86M 98.1±0.1 85.5±0.3 87.2±0.8
DINOv2 ViT-L/14 (f) 300M 99.2±0.0 89.8±0.2 90.1±0.8
PAWS-VMK is competitive with previous semi-supervised methods. Table 1 shows that across a range of benchmarking datasets PAWS-VMK outperforms PAWS using the same frozen backbone model, and is competitive with other methods in the literature that train models of similar sizes and using the same number (or greater) of labelled images. PAWS-VMK outperforms FreeMatch [19] on CIFAR-10 by 0.8% on average with the smallest DINOv2 model available, and also outperform Semi-ViT [18] by 5.1% using a much smaller labelled set. SemiReward [18] obtains better results with a similar model size for CIFAR-100, but in contrast to our approach they fit all model parameters whereas PAWS-VMK uses a frozen backbone. PAWS-VMK outperforms SemiReward (+1.1%) with the ViT-B model, which takes about four hours of GPU time using Nvidia V100 GPUs, compared to the 67 hours of compute time for SemiReward on a more powerful Nvidia A100 GPU [18].
Table 2: Out of distribution detection results on the OpenOOD benchmark [7]. This benchmark reports the average AUROC across the Near-OOD and Far-OOD datasets. Results for PAWS and PAWS-VMK are reported using a frozen DINOv2 ViT-S/14 backbone model. The best performing models are bolded.
Methods Datasets
CIFAR-10 CIFAR-100
Near-OOD Far-OOD Near-OOD Far-OOD
Supervised DeepEnsemble [24, 7] 90.6 93.2 82.7 80.7
KNN [23, 7] 90.5 92.8 79.9 82.2
CIDER [9, 10] 90.7 94.7 72.1 80.5
LogitNorm [25, 7] 92.5 96.7 78.4 81.3
PixMix [51, 7] 93.1 95.7 79.6 85.5
PALM [10] 93.0 98.1 76.9 93.0
Semi-supervised PAWS [8] 82.6 90.5 90.6 94.5
PAWS-VMK 93.1 98.0 95.2 96.3
PAWS-VMK is competitive with previous OOD detection methods. Table 2 shows that PAWS-VMK outperforms PAWS at OOD detection by a large margin. PAWS-VMK also performs similarly to specialised OOD detection methods for the CIFAR-10 OpenOOD [7] benchmark, and outperforms them by a large margin for the CIFAR-100 benchmark. We use the maximum cosine similarity to a labelled prototype as the OOD metric for the PAWS and PAWS-VMK embeddings. One the strengths of our approach is that the in-distribution embeddings are designed to cluster around a small set of prototypes, which makes OOD evaluation much more efficient than other metrics such as KNN [23], which uses a larger sample of the training dataset to determine if a new point is OOD. Other prototypical OOD approaches such as CIDER [9] and PALM [10] employ the KNN or Mahabolonis distance [52] metrics to make predictions, of which the latter typically uses class information which would be unavailable in a semi-supervised setting. The PAWS-VMK projection head also provides more reliable OOD detection than the DINOv2 backbone embeddings (Table S7), and more detailed results for each dataset are provided in the supplementary materials.
Table 3: Out of distribution detection results for different combinations of components on the OpenOOD benchmark [7]. This benchmark reports the average AUROC across the Near-OOD and Far-OOD datasets. The best performing models are bolded.
Methods Components Datasets
CIFAR-10
Head init. Loss type Label selection Strat. Near-OOD Far-OOD
PAWS [8] Random init. Consistency Random Strat. 82.6 90.5
PAWS-M Random init. PAWS-MixMatch Random Strat. 91.4 95.6
PAWS-VM vMF-SNE PAWS-MixMatch Random Strat. 93.2 96.6
PAWS-VMK vMF-SNE PAWS-MixMatch SKMPS 93.1 98.0
Each component of PAWS-VMK improves OOD detection performance. Table 3 shows that PAWS-MixMatch loss, vMF-SNE initialisation and SKMPS all contribute to improving OOD performance over the PAWS baseline. By comparing the OOD detection results for each component being introduced individually, it can be seen that the PAWS-MixMatch loss and vMF-SNE initialisation improve on the Near-OOD and Far-OOD benchmarks over PAWS, and that introducing the SKMPS label selection strategy results in a similar Near-OOD result but better Far-OOD detection.
Table 4: Performance of different prototype selection strategies. The best performing prototype selection strategy for each dataset is bolded.
Method Components Datasets
Head init. Loss type Label selection Strat. C10 C100 Food
PAWS Random init. Consistency Random Strat. 92.9±2.1 70.9±1.5 75.1±2.0
PAWS-K Random init. Consistency SKMPS 94.4±0.5 74.8±0.3 80.5±0.6
Random init. Consistency USL [29] 93.6±0.4 74.4±0.6 79.1±1.1
PAWS-VM vmf-SNE PAWS-MixMatch Random Strat. 95.8±0.1 75.7±0.9 75.9±1.8
PAWS-VMK vmf-SNE PAWS-MixMatch SKMPS 95.9±0.0 77.3±0.4 81.7±0.7
vmf-SNE PAWS-MixMatch USL [29] 95.8±0.1 77.5±0.4 80.0±1.0
Simple $k$-means prototype selection (SKMPS) obtains competitive performance to USL. Table 4 shows that using prototype selection strategies improves the performance of both PAWS and PAWS-VM, and that the best performing approach across both methods and all three datasets considered is our SKMPS strategy. While the results between USL [29] and SKMPS are similar for CIFAR-100, significantly better results are obtained on the CIFAR-10 dataset using PAWS, and on the Food-101 dataset using both PAWS and PAWS-VM. SKMPS also provides more consistent results than USL for the Food-101 dataset, with a smaller variance in performance.
Table 5: Performance improvements from vMF-SNE initialisation and the PAWS-MixMatch loss across a range of datasets. The best performing models are bolded.
Method Components Datasets
Head init. Loss type Label selection C10 C100 EuroSAT DeepWeeds Flowers Food Pets
PAWS[8] Random init. Consistency Random stratified 92.9±2.1 70.9±1.5 94.4±0.3 72.6±4.5 98.5±0.7 75.1±2.0 90.2±1.0
PAWS-VM vMF-SNE PAWS-MixMatch Random stratified 95.8±0.1 75.7±0.9 94.2±0.1 87.8±1.5 98.9±0.6 75.9±1.8 91.6±0.2
vMF-SNE initialisation and the PAWS-MixMatch loss provide performance benefits in combination. In Table 5 we explore the performance of the PAWS-MixMatch loss and vMF-SNE initialisation across seven benchmarking datasets. Significantly better performance is observed in four of these, including CIFAR-10, CIFAR-100, DeepWeeds and Oxford Pets. We conduct an extensive ablation study in Appendix B in the supporting information, and find that for some datasets such as DeepWeeds vMF-SNE initialisation is more significant, while for others such as CIFAR-10 the PAWS-MixMatch loss leads to a larger performance improvement.
5 Discussion
Refer to caption Refer to caption
(a) PAWS
Refer to caption Refer to caption
(b) PAWS-VMK
Figure 2: $t$-SNE visualisation of the projection head embeddings and density plots of minimum cosine distance to a prototype ($\circ$) for in-distribution CIFAR-10 (blue) and OOD CIFAR-100 (orange).
PAWS-VMK learns compact representations. Fig. 2 shows that PAWS-VMK learns a much more compact representation around the labelled prototypes in comparison to PAWS, which has been noted by the authors of CIDER [9] and PALM [10] to improve OOD detection performance. While these methods use a specialised optimization loss that is designed to directly achieve this goal, we show that it can also be achieved by careful alterations to the PAWS framework. Using the PAWS-MixMatch rather than consistency based loss provides a more consistent supervision signal, as the same soft pseudo-label is applied to both augmented views of an image, encouraging more compact embeddings. The vMF-SNE initialisation encourages in-distribution points in the projection head embeddings to be close to other points, through the perplexity parameter described in Eq. 14 in the supporting information. Finally, selecting a representative set of labelled prototypes using SKMPS results in them being more centralised within the in-distribution data in the final embeddings, as can be seen in Fig. 2.
$\Leftarrow$ $\Rightarrow$ $\Rightarrow$
(a) PAWS
(b) Random Init.
(c) vMF-SNE
(d) PAWS-VM
Figure 3: $t$-SNE visualisation of the projection head embeddings at various stages of the training process for DeepWeeds. Blue is the weed present class, and orange is the negative class. The arrows represent the order of training steps.
vMF-SNE initialisation improves the separability of complex class representations. Fig. S8 shows that vMF-SNE initialisation helps PAWS-VMK learn better models when the backbone embeddings clusters them poorly in the first instance. Fig. 3(b) and Fig. 3(c) reflect low dimensional representations of the embeddings of the DINOv2 foundation model backbone on the DeepWeeds dataset, where separation of the weeds versus not weeds classes is quite poor. While training using PAWS results in mediocre performance and a mixing of the two classes (Fig. 3(a)), training using the vMF-SNE initialised head results in better seperation (Fig. 3(d)). We show in Table S8 in the supplementary material that this also holds when vMF-SNE initialisation has been done with a different dataset.
(a) CIFAR-10
(b) CIFAR-100
(c) Food-101
Figure 4: $t$-SNE visualisation of the DINOv2 ViT-S/14 backbone embeddings overlayed with the prototypes selected by the USL ($+$) and SKMPS ($\circ$) strategies.
SKMPS selects more diversity in comparison to USL. Fig. 4 shows that SKMPS samples a more diverse set of prototypes in comparison to USL. We explore this further in the supplementary material, showing in Fig. S8 that SKMPS samples more classes than USL on small budgets. For example, for 202 prototypes on the Food-101 dataset, USL never selects all of the classes in 20 runs whereas SKMPS samples all the classes in 25% of cases. The USL approach considers density to be an indicator of prototype informativeness, and preferentially chooses prototypes from denser regions of the representation space $Z$ [29]. In contrast, SKMPS only considers the density of the points through the $k$-means clustering process, which results in a more diverse sample in addition to being simpler and much more computationally efficient.
6 Conclusion
In this paper we introduced PAWS-VMK, a semi-supervised learning approach based on the PAWS [8] framework that simultaneously achieves state-of-the-art results in semi-supervised learning and OOD detection. PAWS-VMK is designed to be used with foundation models, employing parametric vMF-SNE initialisation and simple $k$-means prototype selection to obtain the maximal benefit from the high-quality embeddings that these models make available. We also find that the original consistency loss proposed by PAWS leads to suboptimal SSL and OOD performance with a frozen foundation model backbone, and instead using the PAWS-MixMatch loss proposed in this paper significantly improves results.
Acknowledgements
We acknowledge the Traditional Owners of the unceded land on which the research detailed in this paper was undertaken: the Wurundjeri Woi-wurrung and Bunurong peoples. This research was undertaken using the LIEF HPC-GPGPU Facility hosted at the University of Melbourne. This Facility was established with the assistance of LIEF Grant LE170100200. This research was also undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government. Evelyn Mannix was supported by a Australian Government Research Training Program Scholarship to complete this work.
Supplementary Material
Appendix A Parametric vMF-SNE and PAWS
The PAWS Similarity kernel and predictions. Following [34], PAWS can be written down as a generative model. The probability density of a sample from the dataset is defined using the prototypes from the labelled dataset $D_{l}$ and a similarity kernel,
$p(x_{j}|D_{l})$ $:=\frac{1}{|D_{l}|}\sum_{x_{i}\in D_{l}}k_{\hat{Z}}(x_{j};x_{i})$ (6)
where the similarity kernel is defined by
$k_{\hat{Z}}(x_{j};x_{i})$ $:=C_{z}\text{exp}\left(\hat{z}_{j}\cdot\hat{z}_{i}/\tau\right)$ (7)
$\hat{z}$ $:=\frac{g_{\phi}(z)}{||g_{\phi}(z)||},\text{ }z:=f_{\theta}(x)$ (8)
with $C_{z}$ a normalizing constant, $f_{\theta}$ is a backbone model, and $g_{\phi}$ is the projection head. PAWS defines this conditional probability within the second latent space $\hat{Z}$, employing the projection head $g_{\phi}$.
For a particular class $y$, we can use the set of prototypes that belong to a class, $B_{y}=\{\{x_{i},y_{i}\}\in D_{l}|y_{i}=y\}$, to define the support of an image conditional on the class $y$
$p(x_{j}|y,D_{l}):$ $=\frac{1}{|B_{y}|}\sum_{x_{i}\in B_{y}}k_{\hat{Z}}(x_{j};x_{i})$ (9)
and through Bayes’ rule, we can derive the predicted class probabilities as
$p(y|x_{j},D_{l})$ $=\frac{p(x_{j}|y,D_{l})p(y)}{p(x_{j}|D_{l})}$ (10)
$=\sigma\left(\hat{z}_{j}\cdot\hat{z}_{S}^{T}/\tau\right)y_{S}=p_{j}$ (11)
where $y_{S}$ is the one-hot encoded or soft labels corresponding to the classes of the prototypes, $\tau$ is a temperature hyperparameter and $\sigma$ is the softmax function. Finally, $z_{S}$ is the matrix of latent representations of the prototypes. The last step assumes class balance in the data and prototypes, setting $p(y)=\frac{1}{C}$ and $B_{y}=N/C$ where $C$ is the total number of classes in the dataset.
Initialising the projection head with vMF-SNE. In this section we describe our parametric Stochastic Neighbour Embedding (SNE) approach [26]. This method involves building two probability distributions that describe the similarity of data points in two different spaces, the origin space $P$ and the target space $Q$. The idea is for similar points $i,j$ in $P$ to have a high probability $p_{ij}$ compared to dissimilar points, and to fit the distribution in the target space $Q$ to the distribution in the origin space $P$ by minimizing the KL divergence
$\text{KL}(Q||P)=\sum_{ij}p_{ij}\log\frac{p_{ij}}{q_{ij}}$ (12)
where $p_{ij}$ and $q_{ij}$ describe the similarity between points in the origin and target space.
These $P,Q$ distributions are built using a probability distribution as a kernel. Taking inspiration from the probablistic perspective of PAWS [34], we propose to use a von Mises-Fisher (vMF) distribution [35] in the origin and target spaces. This is the same as the PAWS similarity kernel in Eq. 7 with the normalizing constant given by $C_{z}=C_{p}(1/\tau)$, with $1/\tau$ the concentration parameter. The normalization constant $C_{p}$ is expensive to compute, but can be efficiently approximated [36].
Using the vMF kernel to compute the probabilities $P$ in the latent space defined by our backbone model $f_{\theta}$, we have
$p_{j|i}$ :=$\frac{k_{Z}(x_{j};x_{i})}{\sum_{k\not=i}k_{Z}(x_{k};x_{i})},$ (13)
$k_{Z}(x_{j};x_{i})$ $:=C_{p}(\kappa_{i})\text{exp}\left(z_{j}^{\prime}\cdot z_{i}^{\prime}\kappa_{i}\right),$ (14)
$z^{\prime}$ $:=\frac{f_{\theta}(x)}{||f_{\theta}(x)||}$ (15)
where we define $p_{i|i}=0$, and note that the normalization constant cancels out in this expression. We set $\kappa_{i}$ for each point to make the entropy $H_{i}$ equal to the perplexity parameter $\gamma$, $H_{i}=\sum_{k}p_{k|i}\log p_{k|i}=\log\gamma$. This ensures each point has a similar number of neighbours in the distribution $P$, and for each point $\kappa_{i}$ is determined using the bisection method. To obtain a symmetric distribution, we use $p_{ij}=\frac{p_{j|i}+p_{i|j}}{2N}$, where $N$ is the number of data points.
To compute the probability distribution $Q$ of the second latent space $\hat{Z}$, a vMF distribution with fixed concentration is used as the kernel,
$q_{j|i}$ =$\frac{k_{\hat{Z}}(x_{j};x_{i})}{\sum_{k\not=i}k_{\hat{Z}}(x_{k};x_{i})}$ (16)
where $k_{\hat{Z}}$ is the PAWS kernel (Eq. 7) and we define $q_{i|j}=0$. We also symmetrise this conditional distribution as above to obtain $q_{ij}$.
To minimize the KL divergence in Eq. 12, gradients are taken with respect to the parameters of the projection head $g_{\phi}$. This allows the vMF-SNE initialisation to follow a similar approach as to training the neural networks in PAWS, utilising mini-batching, stochastic gradient descent and image augmentations [8].
This parametric vMF-SNE initialisation method is a similar to $t$-SNE [28], parametric $t$-SNE [53] and non-parametric vMF-SNE [54], but we apply it in a very different context — as an initialisation approach for neural networks using frozen foundation model embeddings. There are three key hyper-parameters in this process — the batch size, which defines the number of images to sample in each mini-batch, the perplexity parameter $\gamma$, and the concentration of the vMF distributions in $Z_{2}$ which is determined by $\tau$. An ablation study considering the impact of these parameters is presented Section D.2.
Algorithm 1 vMF-SNE initialisation
[…]
Randomly initialise MLP head $g_{\phi}$ and set starting value of $\kappa_{i}=1$ for all images $i$ ;
$l=0$;
while $l<N_{E}$ do
     Randomly split $T$ into $B$ mini-batches;
     for $(x_{b},y_{b})\in\{T_{1},...,T_{b},...,T_{B}\}$ do
         $X=\text{Aug}(x_{b})$
         Calculate $Z^{\prime},Z^{\prime}_{i:}=||f_{\theta}(X_{i})||$;
         Calculate $\hat{Z},\hat{Z}_{i:}=||g_{\phi}(f_{\theta}(X_{i:}))||$;
         Calculate $k_{Z}(x_{j};x_{i})=\text{exp}\left(Z^{\prime}_{j:}\cdot Z^{\prime}_{i:}\kappa_{i}\right)$ for all images $i,j$;
         Calculate $k_{\hat{Z}}(x_{j};x_{i})=\text{exp}\left(\hat{Z}_{j:}\cdot\hat{Z}_{i:}/\tau\right)$ for all images $i,j$;
         $p_{j|i}=\frac{k_{Z}(x_{j};x_{i})}{\sum_{k\not=i}k_{Z}(x_{k};x_{i})}$; $\triangleright$ Normalising constants are divided out.
         $q_{j|i}=\frac{k_{\hat{Z}}(x_{j};x_{i})}{\sum_{k\not=i}k_{\hat{Z}}(x_{k};x_{i})}$;
         Find new values for $\kappa_{i}$ by minimising $\left|\sum_{k}p_{k|i}\log p_{k|i}-\log\gamma\right|$ using bisection method;
         Recalculate $p_{j|i}$ with new values of $\kappa_{i}$;
         Calculate $p_{ij}=\frac{p_{j|i}+p_{i|j}}{2N}$ and $q_{ij}=\frac{q_{j|i}+q_{i|j}}{2N}$;
         Calculate KL divergence $\mathcal{L}=\sum_{ij}p_{ij}\log\frac{p_{ij}}{q_{ij}}$;
         Minimise loss $\mathcal{L}$ by updating $\phi$;
     end for
     $l=l+1$;
end while
Algorithm 2 SKMPS prototype selection
Input: Array of input images $X$, $f_{\theta}$ backbone model, labelling budget $k$
Calculate embeddings of all images in the latent space $Z=f_{\theta}(X)$;
Normalize embeddings to create $K,K_{i:}=||Z_{i:}||$;
Cluster $K$ into $k$ clusters using $k$-means to find the cluster centroids $C$;
Initialise empty list $L=\{\}$;
$i=0$;
while $i<k$ do
     $j=\text{ArgMin}(\text{cosine distance}(K,C_{i:}))$;
     $L=\{L,j\}$;
     $i=i+1$;
end while
Return: List $L$ of the indices of the images to be labelled
Appendix B SSL Ablation studies
Figure S5: Comparison of the vMF-SNE initialisation and PAWS-MixMatch loss components of PAWS-VMK versus the original PAWS approach.
Fig. S5 shows that across the CIFAR-10, CIFAR-100, Food-101 and DeepWeeds datasets, these components together lead to the best overall results. For CIFAR-10 and CIFAR-100, the PAWS-Mixmatch loss significantly improves overall accuracy and the gain from vMF-SNE initialisation is less obvious. However, for DeepWeeds skipping vMF-SNE initialisation leads to poor results. In the case of the Food-101 dataset, the combination of both vMF-SNE initialisation and the PAWS-MixMatch loss leads to be best average accuracy across all training runs.
Appendix C Additional results
Figure S6: Impact of sharpening approach in PAWS on the class probability. The horizontal axis describes probability of the first class, and vertical axis describes the sharpened probability when $T=0.25$ in Equation 3, with the probability of all other classes set to be equal in the input.
Figure S7: Training curves comparing the vMF-SNE and random initialisation of the projection head and consistency versus the PAWS-Mixmatch loss.
Consistency based loss can degrade the performance of PAWS during a training run. In Fig. S7 the training curves for two of the datasets from Fig. S5 are shown. These training curves demonstrate that the consistency loss can result in performance degrading for CIFAR-10 after around 10 epochs. When using the PAWS-Mixmatch loss, the validation accuracy remains stable once the models have converged. It is also noted that the performance improvement from vMF-SNE initialisation is not due to a head-start in training. While it does provide a better starting point for learning, in both cases models are trained until they completely converge, and vMF-SNE initialised projection heads converge to higher performing maxima, particularly for the DeepWeeds dataset where training using a randomly initialised projection head leads to poor performance.
(a) Average number of classes sampled over 20 runs. Error bars show maximum and minimum number of classes sampled.
(b) Proportion of sets sampling all classes over 20 runs.
Figure S8: Class diversity captured by different prototype selection strategies.
SKMPS selects a more diverse sample than USL. In Fig. S8 we compare diversity of classes sampled for the USL and SKMPS strategies for progressively larger budgets for the CIFAR-10, CIFAR-100 and Food-101 datasets. In all three cases, the SKMPS strategy samples a greater diversity of images, as measured by the number of classes sampled from each dataset for a given budget. When not all classes are sampled, this reduces the maximum attainable performance for a model as classes that are not represented by atleast one labelled image cannot be learned in most semi-supervised learning frameworks.
Table S6: Detection performance on near OOD datasets with the DINOv2 ViT-S/14 frozen backbone. The $\uparrow$ means larger values are better and the $\downarrow$ means smaller values are better. Bolded numbers indicate the best result for a particular dataset. The left-most column indicates the in-distribution-dataset (IDD).
IDD Method Near OOD Datasets Average
CIFAR-10 CIFAR-100 Tiny ImageNet
[…]
CIFAR-100 PAWS 90.11 51.13 90.99 46.54 90.55 48.84
PAWS-VMK 95.79 26.05 94.60 32.30 95.19 29.18
backbone 85.51 60.16 91.13 30.6 88.32 45.38
CIFAR-10 PAWS 82.10 73.90 83.19 69.49 82.64 71.70
+PAWS-MixMatch 91.61 43.36 91.12 42.78 91.37 43.07
+vMF-SNE 92.84 40.43 93.50 34.92 93.17 37.68
PAWS-VMK 93.34 35.94 92.83 33.30 93.09 34.62
backbone 90.66 40.13 97.16 11.12 93.91 25.62
Table S7: Detection performance on far OOD datasets with the DINOv2 ViT-S/14 frozen backbone. The $\uparrow$ means larger values are better and the $\downarrow$ means smaller values are better. Bolded numbers indicate the best result for a particular dataset. The left-most column indicates the in-distribution-dataset (IDD).
IDD Method Far OOD Datasets Average
DTD MNIST SVHN Places365
[…]
CIFAR-100 PAWS 93.59 35.71 97.30 13.95 94.33 32.66 92.78 39.65 94.50 30.49
PAWS-VMK 96.55 22.29 98.96 2.27 94.35 33.39 95.22 28.58 96.27 21.63
backbone 98.92 4.8 87.98 86.15 72.76 85.57 97.03 12.18 89.17 47.18
CIFAR-10 PAWS 91.36 48.71 96.44 15.37 84.99 81.74 89.27 54.24 90.52 50.01
+PAWS-Mixmatch 95.44 27.84 98.67 5.35 93.81 39.49 94.48 27.4 95.6 25.02
+vMF-SNE 96.05 24.95 98.68 1.91 95.50 28.77 96.06 22.20 96.57 19.46
PAWS-VMK 97.45 15.62 99.78 0.00 98.04 10.78 96.70 17.13 97.99 10.88
backbone 99.96 0.16 99.4 1.91 80.57 71.18 98.8 4.17 94.68 19.36
The PAWS-VMK projection head provides more reliable OOD detection than the backbone embedding. In Table S6 and Table S7 we present the full set of results for each dataset in the OpenOOD benchmark. We also provide results for the backbone embedding and compare them to the PAWS and PAWS-VMK projection heads. For the backbone approach, we select prototypes using the SKMPS strategy and use the smallest cosine distance to a prototype as our OOD metric, as done for PAWS-VMK. In general we find that the backbone embedding performs better on the datasets that are included in the DINOv2 training set, such as ImageNet, DTD and Places365 (which is included in the ADE20K dataset [55]). However, the backbone performs very poorly on OOD detection for datasets it hasn’t seen before such as CIFAR-10, CIFAR-100, MNIST and SVHN, while our projection heads, particularly the PAWS-VMK embeddings, can still perform well. Overall, the backbone performs slightly better on the near OOD CIFAR-10 benchmark (+0.82), while the PAWS-VMK embeddings perform better on the far OOD CIFAR-10 (+3.31) and both CIFAR-100 (+6.87/+7.10) benchmarks.
Table S8: Transferability of vMF-SNE initialisation between datasets.
Method Components Dataset
Head init. Loss type Label selection DeepWeeds
Random init. PAWS-Mixmatch Random stratified 76.0±9.3
vMF-SNE init. on C100 PAWS-Mixmatch Random stratified 79.9±8.1
vMF-SNE init. on C10 PAWS-Mixmatch Random stratified 81.5±7.1
vMF-SNE init. on Food PAWS-Mixmatch Random stratified 84.2±6.0
PAWS-VM vMF-SNE init. on DeepWeeds PAWS-Mixmatch Random stratified 87.8±1.5
Projection heads initialised using the vMF-SNE approach perform better than random initialisation, even if they are initialised using a different dataset. In (Table S8) we fit models on the DeepWeeds dataset randomly initialising the projection head, and then use projection heads initialised on various other datasets using our vMF-SNE approach. We found that initialising the projection head with vMF-SNE is better than random initialisation in every case, but best performance was obtained when initialisation was undertaken on DeepWeeds. Potentially, the vMF-SNE approach could be used as a method to initialise a projection head for a foundational model on a large dataset, and this could then be used to obtain better performance than random initialisation on smaller datasets.
Table S9: Compatability of a MixMatch inspired loss and vMF-SNE initialisation with RoPAWS.
Method Components Dataset
Head init. Loss type Label selection C10 DeepWeeds
RoPAWS [34] Random init. RoPAWS-Consistency Random stratified 93.3±1.5 81.2±6.8
RoPAWS-VM vMF-SNE RoPAWS-Mixmatch Random stratified 95.7±0.1 86.4±0.6
vMF-SNE initialisation and a Mixmatch inspired loss also provide performance improvements when used with RoPAWS. It is shown in Table S9 that the improvements to the PAWS framework for learning with a frozen backbone model also improves the performance obtained when using RoPAWS [34] on in-distribution data. This suggests that the characteristics of PAWS that leads to poor performance on DeepWeeds and CIFAR-10 also hold for RoPAWS.
Table S10: Adding terms from the PALM [10] loss that improve OOD detection performance to PAWS [8] and PAWS-M.
Method Components Dataset
Loss type PALM [10] $\mathcal{L}_{\text{proto-contra}}$ C10
PAWS Consistency Y 92.1±0.1
PAWS Consistency N 92.9±0.2
PAWS-M PAWS-Mixmatch Y 94.2±0.1
PAWS-M PAWS-Mixmatch N 95.9±0.1
Naively incoporating elements of the PALM loss into PAWS harms SSL performance. It is shown in Table S10 that incorporating the loss term from PALM that encourages prototypes within a class to be compact, and far from prototypes of other classes, given by
$\mathcal{L}_{\text{proto-contra}}=-\frac{1}{CK}\sum_{c=1}^{C}\sum_{k=1}^{K}\log\frac{\sum_{k^{\prime}=1}^{K}\mathbbm{1}(k^{\prime}\not=k)\exp{(\hat{z}_{Sc}^{k^{\prime}}\hat{z}_{Sc}^{k\top}/\tau)}}{\sum_{c^{\prime}=1}^{C}\sum_{k^{\prime\prime}=1}^{K}\mathbbm{1}(k^{\prime\prime}\not=k,c^{\prime}\not=c)\exp{(\hat{z}_{Sc^{\prime}}^{k^{\prime\prime}}\hat{z}_{Sc^{\prime}}^{k\top}/\tau)}}$ (17)
results in poorer performance in comparison to not adding this loss term. Here $\mathbbm{1}$ is an indicator function that avoids contrasting between the same prototype.
Appendix D Additional results for vMF-SNE initialisation
D.1 Performance
Table S11: KNN accuracy of the model backbone versus the vMF-SNE initialised projection head. Bolded entries indicate the best results for a particular dataset.
Embedding Dataset
C10 C100 EuroSAT DeepWeeds Flowers Food Pets
DINOv2 ViT-S/14 backbone 96.2 82.3 90.0 88.7 99.3 80.9 91.5
vMF-SNE proj. head 96.3±0.1 79.0±0.1 91.9±0.1 91.3±0.1 98.2±0.5 81.0±0.1 93.0±0.1
In a classification context, the performance of the backbone model can be measured non-parametrically using a weighted k-Nearest Neighbor (kNN) Classifier [56]. This provides for a quantitative approach to measure the performance of the projection head initialised using vMF-SNE, as reproducing the local structure of the backbone model should provide a similar kNN accuracy. We find that for five of the seven datasets considered, we can meet or beat the kNN accuracy of the backbone model through vMF-SNE initialisation (Table S11).
D.2 Hyperparameter ablation study
Table S12: Ablation study on the vMF-SNE initialisation hyperparameters. The results for the parameters used in the other results in this paper are reported as the mean KNN accuracy and standard deviation of five runs, while the other cases report a single run.
Perplexity ($\gamma$) C10 C100 Food
5 96.1 74.0 81.7
30 96.3±0.1 79.0±0.1 81.0±0.1
50 96.3 79.5 80.9
100 96.3 80.1 80.7
vMF concentration ($\tau$) C10 C100 Food
0.01 96.4 79.4 78.8
0.10 96.3±0.1 79.0±0.1 81.0±0.1
1.00 91.7 54.9 68.5
$Z_{2}$ dimension C10 C100 Food
128 96.3 79.1 81.0
256 96.3 79.0 81.0
384 96.4 79.1 81.0
512 96.3±0.1 79.0±0.1 81.0±0.1
1024 96.3 79.0 81.0
Batch size C10 C100 Food
128 96.2 79.5 80.4
256 96.3 79.1 80.7
512 96.3±0.1 79.0±0.1 81.0±0.1
1024 96.3 79.2 81.3
Kernel C10 C100 Food
vMF 96.3±0.1
Gaussian 91.4±0.6
We undertook an ablation study to determine the sensitivity of vMF-SNE initialisation performance to the hyper-parameters within the algorithm. It was found that the method was highly insensitive to most parameters for the CIFAR-10 dataset, with only a significantly lower perplexity $\gamma$ or an order of magnitude higher concentration parameter $\tau$ resulting in changes to the kNN validation accuracy outside a single standard deviation (Table S12). The optimal parameter set appeared to be different between CIFAR-100 and the Food-101 datasets, with better performance for CIFAR-100 with a larger perplexity and smaller concentration and batch size, but better performance for Food-101 with a smaller perplexity and larger batch size. We also explored using a Gaussian kernel with CIFAR-10, and found that this resulted in much poorer performance.
Appendix E Further implementation details
E.1 PAWS
We adapted the PAWS and RoPAWS PyTorch [57] implementation available at this GitHub repository to our own codebase using PyTorch-Lightning [58] and Hydra [59], which we ensured produced identical results. We employed the same samplers, optimizers, scheduler, augmentations, projection head architecture et cetera as described in [8]. The hyperparameters we selected are described in Appendix F.
E.2 vMF-SNE initialisation
We re-used the same augmentations, optimizer, scheduler and other details as for PAWS. The key difference is that the vMF-SNE initialisation approach only requires the unlabelled dataloader, compared to PAWS which requires both a labelled and unlabelled dataloader. This meant vMF-SNE initialisation was slightly faster. We note than when using image augmentations for this approach, we treat each augmented view independently.
E.3 $k$-means clustering and USL
We used the gpu-optimised $k$-means method available within cuML [60]. We save the normalised global class token from the DINOv2 model applied to each image of the dataset using the validation transformations, and run the $k$-means clustering method over this matrix. This method is extremely fast and can easily process millions of rows. We use the best clustering result from ten runs, and choose the number of clusters according to the number of images we wish to label. To select the prototype to label from each cluster, we choose the image with the largest cosine similarity to the cluster centroid.
To reproduce the USL [29] approach, we used the code published by the authors on GitHub. There are quite a few hyperparameters that are used within the USL method, so we selected the set that was used for ImageNet with CLIP [12], another foundation model based on the ViT architecture, for all of our experiments.
Appendix F Hyperparameters and further details
F.1 PAWS
We closely follow the hyperparameters suggested for PAWS for RoPAWS and use the same set for every dataset. These are summarised in the table below.
[…]
Hyperparameter Value
[…]
Hyperparameter Value
[…]
{xtabular}ll Sampler parameters &
Labelled Sampler Class-stratified
Labelled batchsize 320
Classes per batch All*
Unlabelled Sampler Random
Unlabelled batchsize 512
Optimiser parameters
Optimizer LARS SGD
Weight decay 1e-6
Momentum 0.9
Scheduler parameters
LR schedule Cos. anneal. w/ lin. warmup
Start LR 0.3
Maximum LR 6.4
Final LR 0.064
Epochs 50
Training augmentations
Global crop size 224
Global crop scale (0.14, 1.0)
Local crop size 98
Local crop scale (0.05, 0.14)
Color distortion $s$ 0.5
Blur probability 0.0
Color distortion (grayscale) No
Color distortion (solarize) Yes
Color distortion (equalize) Yes
normalize (mean) (0.485, 0.456, 0.406)
normalize (SD) (0.229, 0.224, 0.225)
Training augmentations Labelled
Global crops 2
Local crops 0
Training augmentations Unlabelled
Global crops 2
Local crops 6
Validation augmentations
Resize 256
Center crop 224
normalize (mean) (0.485, 0.456, 0.406)
normalize (SD) (0.229, 0.224, 0.225)
PAWS
Label smoothing 0.1
Sharpening $T$ 0.25
Temperature 0.1
Me-max Yes
RoPAWS
Prior temperature 0.1
Prior power 1.0
Label ratio 5.0
S-batchsize 32
U-batchsize 512
Projection head
Input dimension 384**
Hidden dimension 384
Output dimension 512
Layers 3
*Except for Food-101 where 100 classes per batch were selected. **This is the size of the output global token from the DINOv2 ViT-S/14 distilled model.
F.2 vMF-SNE initialisation
We use the same parameters as above, with the following additions and changes.
[…]
Hyperparameter Value
[…]
Hyperparameter Value
[…]
{xtabular}ll Sampler parameters
Labelled Sampler None
Labelled batchsize -
Unlabelled Sampler Random
Unlabelled batchsize 512
vMF-SNE parameters
vMF concentration ($\tau$) 0.1
Perplexity ($\gamma$) 30
kNN evaluation parameters
Nearest neighbours 200
Temperature 0.1