### v4

# A Unified Approach to Semi-Supervised Out-of-Distribution Detection

Abstract
One of the early weaknesses identified in deep neural networks trained for image classification tasks, was their inability to provide low confidence predictions on out-of-distribution (OOD) data, that was significantly different from the in-distribution (ID) data used to train them. Representation learning, where neural networks are trained in specific ways that improve their ability to detect OOD examples, has emerged as a promising direction to solving this problem. However, these approaches require long training times, and can be computationally inefficient at detecting OOD examples. Recent developments in Vision Transformer (ViT) foundation models—large networks trained on large and diverse datasets with self-supervised approaches—also show strong performance in OOD detection, and could potentially address some of these challenges. This paper presents Mixture of Exemplars (MoLAR), an approach that provides a unified way of tackling OOD detection challenges in both supervised and semi-supervised settings—that is designed to be trained with a frozen, pretrained foundation model backbone. MoLAR is efficient to train, and provides strong OOD performance when only comparing the distance of OOD examples to the exemplars, a small set of images chosen to be representative of the dataset. As a result, determining if an image is OOD with MoLAR is no more expensive than classifying an image. Extensive experiments demonstrate the superior OOD detection performance of MoLAR in comparison to comparable approaches, and also the strong performance of MoLAR in semi-supervised settings. Code is available on GitHub.
1 Introduction
(a) Supervised.
(b) Self/Semi-supervised.
Figure 1: Effectiveness of OOD detection with exemplars. Comparison of Mixture of Exemplars (MoLAR) against comparable approaches, trained using a frozen DINOv2 ViT-S/14 backbone. MoLAR outperforms PALM, CIDER and PAWS across most in-distribution—out-of-distribution pairs, and improves on the DINOv2 backbone, particularly for datasets not within the foundation model training set.
From managing weeds [26], to food safety [33], to diagnosing cancer [12], deep learning is used extensively in computer vision applications where identifying out-of-distribution (OOD) data, such as plants or food contaminants not present in the training dataset, can be of great importance. While the original methods of training neural networks from scratch resulted in poor correlations between confidence and OOD data [24], modern foundation models [27] — models trained on a large and diverse dataset using self-supervised learning methods that offer quick training and strong performance even when the backbone is frozen — can achieve sound OOD detection results by comparing the distances between in-distribution (ID) and OOD image embeddings.
While distance-based OOD detection metrics have been well developed in the literature [35], they can be improved by developing an embedding that tightly clusters classes and while also maximizing the distance between them. CIDER [21] was the first approach to demonstrate this, and utilized a von Mises-Fisher (vMF) mixture modelling approach with an additional loss that maximised the distance between clusters. PALM [20] extended upon this approach to enable each class to be modelled by multiple mixture components, and found this resulted in improved OOD detection performance.
While effective, these methods find it challenging to learn the vMF cluster centers and require long training times. In CIDER, the cluster centers are not directly optimized using backpropogation, but instead are learned using an Exponential Moving Average (EMA) approach. PALM also learns the cluster centers using the same approach, and assigns the mixture components to images within a batch using the iterative Sinkhorn-Knopp algorithm, adding additional complexity. Further, both of these approaches avoid using the learned cluster centers for OOD detection, and instead employ the KNN [35] or Mahalanobis [17] approach. This is a missed opportunity in making OOD detection efficient, as if distance to the vMF clusters could be used to determine OOD examples, this would not add an additional computational burden for models to screen OOD data.
It is also unclear whether these approaches will work well with foundation models. They are designed to train a convolutional neural network (CNN) backbone with a projection head—a MLP network with a few layers—and have not been designed with a frozen pretrained Vision Transformer (ViT) backbone, which could significantly reduce training times. There are other prototypical deep learning approaches that could be adapted to OOD detection, that can train vMF cluster centers within a mixture model simply and efficiently. In the PAWS [1] semi-supervised learning approach, cluster centers are defined by the embeddings of a small set of labelled data and models are trained using a consistency based loss. However, we find PAWS to perform poorly in OOD detection settings.
To solve these challenges, this paper presents Mixture of Exemplars (MoLAR), an OOD detection approach that uses exemplars—particular images within the training dataset that define the centers of the vMF components of a mixture model—to obtain state-of-the-art results in OOD detection. MoLAR is able to obtain strong performance (Fig. 1) using only distances from exemplars to determine if a point is OOD, making testing for OOD data free and efficient in comparison to other methods. Further, MoLAR provides a unified way of tackling OOD detection challenges in both supervised and semi-supervised settings, while still achieving competitive performance. In this paper, we:
  • Describe MoLAR, a unified approach to OOD detection that can handle both supervised and semi-supervised settings, designed to be used with a frozen foundation model backbone. Unlike previous comparable approaches such as PALM and CIDER, MoLAR is designed to identify OOD examples in a way that is no more computationally expensive than making a class prediction.
  • Demonstrate MoLAR obtains state-of-the-art OOD detection performance in both supervised and semi-supervised settings with a small amount of labelled data, employing the OpenOOD [46] benchmarks.
  • Show that MoLAR-SS, the first semi-supervised method that obtains similar OOD detection performance in comparison to supervised approaches, obtains competitive accuracy with state-of-the-art semi-supervised learning models. Further, the design of MoLAR contributes to strong semi-supervised accuracy across a broader range of datasets.
2 Related work
Out-of-distribution detection.
The problem of out-of-distribution (OOD) detection in deep learning was introduced when it was observed that neural networks tend to provide overconfident results when making predictions on images with previously unseen classes or that are unrecognizable [24]. OOD detection methods are designed to detect when inference is going to be made on such data, so an appropriate alternative action can be taken [36].
Representational approaches such as PALM and CIDER are designed to learn neural networks that produce an embedding of an image that more accurately identifies OOD images under particular OOD detection metrics. Key metrics include KNN [35], which uses the distance of a point to its $k^{\text{th}}$ nearest neighbour in a sample of the training dataset; and the Mahalanobis distance [17], a measure of distance between a point in high dimensional space and a distribution derived from the training data, that uses class means and a pre-calculated covariance matrix. Other works improve OOD detection in a variety of ways, from using particular types of image augmentations [10], employing ensembles [14] to adding corrections to the logits to better capture OOD examples [43].
This paper focuses on improving representational OOD detection methods where only the ID data is available for training. In this context benchmarking OOD detection approaches can be challenging, as there are a large number of different ID and OOD dataset pairs that can be used. This paper follows the OpenOOD benchmarks [46], to ensure better comparability with other works in this space. Most OOD detection methods assume class labels are available for all in-distribution images, and a few have considered the entirely unsupervised case [20]. We are unaware of any previous OOD detection methods that have been designed for a semi-supervised context.
Semi-supervised learning.
In semi-supervised learning (SSL), a model is trained to perform a particular task, such as image classification, using a small set of labelled data and a much larger set of unlabelled data. By taking into account the distribution of the unlabelled data within the learning process, SSL methods can train more accurate models compared to approaches that only utilize the labelled data within this context [39]. The earliest SSL approaches used pseudolabelling [16], and a number of modern approaches have extended on this idea such as FixMatch [34], which employs image augmentation to efficiently propogate labels. Further methods have improved on FixMatch by better accounting for the geometry of the image embeddings [18, 42, 4]. Of all these approaches, PAWS [1] has striking similarities to CIDER and PALM. All three methods use a vMF mixture modelling approach, though where PAWS uses the labelled data to generate cluster centers CIDER and PALM employ a non-parametric approach. Notably, where CIDER and PALM are designed to learn a compact representation of the training data, PAWS uses a consistency based loss where labels are swapped between different views of an image to provide a semi-supervised learning signal.
Foundation models.
In computer vision, foundation models can be used as functional maps from the image domain to a representation space that captures the semantic content of the image with minimal loss of information. The degree to which semantic content is retained can be tested by undertaking downstream tasks within the representation space, such as image classification, segmentation and object detection [31]. Current state-of-the-art foundation models include methods that combine several self-supervised approaches on large datasets, such as DINOv2 [27], and other approaches that use language-image pretraining approaches such as CLIP [30, 5].
Frozen backbone $f$ Augmented views of image $x,x^{+}$ Projection head $g_{\theta}$ Image embeddings $z,z^{+}$ Exemplars $x_{e}\in D_{e}$ Frozen backbone $f$ Projection head $g_{\theta}$ Exemplar embeddings $\hat{\textbf{z}}_{e}$ Class prediction $p,p^{+}$ Supervised loss $H\left(y,p\right)+H\left(y,p^{+}\right)$ Semi-supervised loss $H\left(\rho(\frac{p+p^{+}}{2}),p\right)$
Figure 2: Training framework for MoLAR. Using a vMF mixture modelling approach within the image embeddings, MoLAR makes classifications on the basis of a set of exemplars with known labels. In the supervised case, a cross-entropy loss is employed. In the semi-supervised case, predictions are averaged and sharpened to provide a strong supervision signal.
Frozen backbone $f$ Augmented views of image $x,x^{+}$ Backbone embeddings $f(x),f(x^{+})$ $P$ Projection head $g_{\theta}$ Image embeddings $z,z^{+}$ vMF-SNE loss $D_{\text{KL}}(P||Q)$ SKMPS exemplar selection $Q$
Figure 3: Initialisation framework for MoLAR. The SKMPS strategy is used to select representative exemplars from the dataset. The vMF-SNE approach is used to initialise the embeddings of the projection head—to reflect the local structure of the high quality foundation model embeddings.
3 Methodology
Motivation.
Foundation models are trained to produce embeddings that capture semantic information across a wide range of datasets. When these models are used for a particular task, a classification head can be trained on the frozen foundation model backbone using a much smaller target dataset, providing strong results with a fraction of the training cost of fine-tuning the model backbone. Further, we find that the model backbone provides a good representation for identifying OOD examples—some of which, were likely seen in the training dataset used for learning the foundation model in the first instance. Nevertheless, this may lead to particular biases in OOD detection when using the foundation model backbone as shown in (Fig. 1), with different performance on OOD data depending on if it was seen during self-supervised pretraining. Nevertheless, foundation models provide a promising direction to solve some of the current challenges with state-of-the-art OOD detection approaches based on learning representations, from the need for long training times in comparison to standard supervised approaches to additional overhead being required to use these approaches to identifying OOD examples.
In this work, we consider if it is possible to get the best of both worlds, and design Mixture of Exemplars (MoLAR) — an OOD detection approach that synergises with a strong foundation model backbone, while providing a highly efficient approach to identifying OOD data that is no more expensive than making a classification.
Notation.
[…]
Supervised MoLAR.
We define MoLAR as a mixture model with von Mises-Fisher components
$p(\hat{z}|D_{e}):$ $=\frac{1}{|D_{e}|}\sum_{x_{i}\in D_{e}}k(\hat{z};\hat{z}_{i})$ (1)
$k(\hat{z}_{j};\hat{z}_{i})$ $\propto\text{exp}\left(\hat{z}_{j}\cdot\hat{z}_{i}/\tau\right)$ (2)
where $z_{i}=g_{\theta}(f(x_{i}))$, inputs are normalized onto the unit sphere $\hat{z}_{j}=\frac{\hat{z}_{j}}{||\hat{z}_{j}||}$ and $\tau$ is a temperature hyperparameter that sets the diffuseness of the distribution. For a particular class $y$, we have
$p(\hat{z}|y,D_{e})$ $=\frac{1}{|D_{e}^{y}|}\sum_{x_{i}\in D_{e}^{y}}k(\hat{z};\hat{z}_{i})$ (3)
where $D_{e}^{y}$ are the exemplars of class $y$. Using Bayes’ rule provides a way to classify new examples $z$ with
$p(y|\hat{z},D_{e})$ $=\frac{p(\hat{z}|y,D_{e})p(y)}{p(\hat{z})}$ (4)
$=\sigma(\hat{z}\cdot\hat{\textbf{z}}_{e}/\tau)\cdot\bm{\phi}$ (5)
where we assume a uniform prior $p(y)$, and use $\hat{\textbf{z}}_{e}$ to refer to the matrix of the normalized exemplar embeddings, while $\bm{\phi}$ is the matrix of their one-hot encoded class labels. This provides for a straightforward maximum likelihood estimation approach for training the model
$\mathcal{L}_{\text{supervised}}(D,D_{e};\theta):=\frac{1}{|D|}\sum_{x_{i},y_{i}\in D}p(y_{i}|\hat{z_{i}},D_{e})$ (6)
which is equivalent to a cross entropy loss.
Semi-Supervised MoLAR.
In the semi-supervised setting, we do not have the labels $y$ for the whole dataset $D$, only our exemplars $D_{e}$. This means an alternate loss function is required, and we find that averaging and sharpening predictions over multiple augmented views $x,x^{+}$, similar to MixMatch [2], provides best results
$\mathcal{L}_{\text{semi-supervised}}:$ $=\frac{1}{2|D|}\sum_{i=1}^{|D|}\left(H\left(\rho\left(\frac{p_{i}+p_{i}^{+}}{2}\right),p_{i}\right)\right.$ (7)
$+\left.H\left(\rho\left(\frac{p_{i}+p_{i}^{+}}{2}\right),p_{i}^{+}\right)\right)-H(\overline{p}),$
$\rho(p)_{i}:$ $=\frac{p_{i}^{1/T}}{\sum_{i=1}^{C}p_{i}^{1/T}}$ (8)
where $p_{i}=p(y|\hat{z_{i}},D_{e})$ as given by Eq. 5, $\rho(\cdot)$ is a sharpening function with temperature $T$, and $H(p,q)=\sum_{i}p_{i}\log q_{i}$ is the cross-entropy term. In the final expression, $\overline{p}=\sum_{i}^{|D|}\rho\left(\frac{p_{i}+p_{i}^{+}}{2}\right)$ is the average of the sharpened predictions on the unlabelled images, which encourages the predictions to be balanced across the classes. We refer to this approach as MoLAR-SS.
Exemplar selection strategies.
In this step, the selection of exemplars $D_{e}$ from the whole dataset $D$ is considered. The goal is to select the set of exemplars, that are most representative of the data as evidenced by better downstream accuracy and OOD detection performance. In the default case, exemplars can be randomly selected uniformly across classes. However, this requires all the data to be labelled in the first instance, which is at odds with a semi-supervised learning context.
In this work, a simple $k$-means exemplar selection (SKMPS) strategy is found to provide good performance. This approach provides for a diverse sampling of exemplars in three steps: (1) calculate the normalized foundation model embeddings of the images, then (2) cluster them into $k$ clusters, where $k$ is the desired number of exemplars (i.e. number of labelled images in a semi-supervised setting). And (3) select the image closest to the centroid of each cluster as an exemplar.
We choose $k$-means clustering [19] for the clustering algorithm in the second step as it is fast to compute and allows for the selection of a fixed number of clusters. The $k$-means algorithm minimises the sum of squared euclidean distances to the cluster centroids [19]. By normalizing the image embeddings given by the foundation model in the first step, the squared euclidean distance becomes proportional to cosine distance, which is used for determining image similarity with foundation models [27, 30]. In this sense, the SKMPS procedure can be interpreted as identifying clusters of visually similar images, and selecting the most representative image from each cluster as an exemplar.
Projection head initialisation.
If a pretrained backbone model $f$ is used, which provides a high quality representation of the dataset, a good starting point for the projection head $g_{\theta}$ should reflect the local structure of the data under the image of $f$. This is not a given if the parameters of the projection head are randomly initialized, and without placing restrictions on its architecture and the dimensionality of the latent space of $f$ and $g_{\theta}$, the projection head cannot be initialized as an identity mapping.
Instead, a Stochastic Neighbour Embedding (SNE) [11] approach is proposed. This involves building two probability distributions $P,Q$ using a kernel that describes the similarity of data points in the origin space (under the image of $f$) and the target space (under the image of $g_{\theta}\circ f$). The idea is for images $x_{i},x_{j}$ in $D$, $p_{ij}\in P$ will have high probability if they are similar and low probability if they are dissimilar. The target distribution $Q$ in the target space can then be fit to the distribution in the origin space $P$ by minimizing the KL divergence
$D_{\text{KL}}(P||Q)=\sum_{ij}p_{ij}\log\frac{p_{ij}}{q_{ij}}$ (9)
where $p_{ij}$ and $q_{ij}$ describe the similarity between points in the origin and target space. Inspired by the definition of MoLAR as a vMF mixture approach, a vMF kernel is used to calculate $P,Q$ as described in Appendix A, providing for an effective method of initializing the parameters of the projection head.
Table 1: OOD detection. Comparison of methods on the OpenOOD benchmark [46]. We use — to refer to the performance of the backbone embeddings (without a projection head)
Backbone Methods Datasets
CIFAR-10 CIFAR-100
Near-OOD Far-OOD Near-OOD Far-OOD
ResNets DeepEnsemble [14, 46] 90.6 93.2 82.7 80.7
(finetuned) KNN [35, 46] 90.5 92.8 79.9 82.2
LogitNorm [43, 46] 92.5 96.7 78.4 81.3
PixMix [10, 46] 93.1 95.7 79.6 85.5
CIDER [21, 20] 90.7 94.7 72.1 80.5
PALM [20] 93.0 98.1 76.9 93.0
DINOv2 ViT-S/14 95.8 98.1 89.9 92.2
(frozen) CIDER 92.4 97.6 76.5 91.2
PALM 94.3 98.6 86.4 95.7
MoLAR 96.8 98.2 95.5 96.9
Comparison to closely related works.
MoLAR is similar to CIDER [35] and PALM [20], but the vMF cluster centers are defined by the embeddings of the examplars in $D_{e}$ and no loss term is added to separate components from different classes. This additional term is not found to be necessary to obtain good results, and hurts accuracy in semi-supervised tasks (Tab. S15). MoLAR is also similar to PAWS [1], but we (1) additionally consider an entirely supervised setting; (2) replace the original consistency based loss with an averaged and sharpened loss; (3) do not use an additional classification head, reporting accuracy only using Eq. 5; and (4) consider exemplar selection strategies and initialise the projection head with the foundation model embeddings, which provides better semi-supervised accuracy and OOD detection performance in this setting. The SKMPS exemplar selection strategy is also similar to Unsupervised Selective Labelling (USL) [41], which is proposed to improve semi-supervised learning performance by selecting a better set of labelled images in the first instance. The key difference between the two approaches is that SKMPS does not require estimating the density of the image embeddings from the dataset, making it simpler to apply and more computationally efficient.
Table 2: OOD detection with exemplars. Comparison of methods on the OpenOOD benchmark [46], using a consistent set of exemplars obtained with SKMPS. We use — to refer to the performance of the backbone embeddings (without a projection head).
Backbone Methods Datasets
CIFAR-10 CIFAR-100
Near-OOD Far-OOD Near-OOD Far-OOD
DINOv2 ViT-S/14 Self-supervised 93.9 94.7 88.3 89.2
(frozen) Supervised CIDER 90.7 96.8 78.6 91.1
PALM 91.7 97.1 85.2 94.5
MoLAR 95.4 97.2 95.9 97.0
Semi-supervised PAWS 82.6 90.5 90.6 94.5
MoLAR-SS 93.1 98.0 95.2 96.3
Table 3: Semi-supervised learning. Performance of MoLAR versus PAWS and other methods. Comparable model sizes in each column are underlined, and the best performing model of these is bolded. Entries noted with * use a larger labelled dataset (1% of the data) than in our benchmarks.
Methods Backbone Datasets
C10 C100 Food
FreeMatch [47] WRN-28-2 1M 95.1
+SemiReward [18] ViT-S-P4-32 21M 84.4
Semi-ViT [4] ViT-Base 86M 82.1*
PAWS DINOv2 ViT-S/14 (f) 21M 92.9±2.1 70.9±1.5 75.1±2.0
MoLAR-SS DINOv2 ViT-S/14 (f) 21M 95.9±0.0 77.3±0.4 81.7±0.7
DINOv2 ViT-B/14 (f) 86M 98.1±0.1 85.5±0.3 87.2±0.8
DINOv2 ViT-L/14 (f) 300M 99.2±0.0 89.8±0.2 90.1±0.8
Table 4: Exemplar selection strategies. Comparison of different exemplar selection strategies with MoLAR-SS and PAWS.
Method Exemplar sel. strat. Datasets
C10 C100 Food
PAWS Random Strat. 92.9±2.1 70.9±1.5 75.1±2.0
USL 93.6±0.4 74.4±0.6 79.1±1.1
SKMPS 94.4±0.5 74.8±0.3 80.5±0.6
MoLAR-SS Random Strat. 95.8±0.1 75.7±0.9 75.9±1.8
USL 95.8±0.1 77.5±0.4 80.0±1.0
SKMPS 95.9±0.0 77.3±0.4 81.7±0.7
Table 5: Semi-supervised learning on additional datasets. Comparison of PAWS and MoLAR-SS on a wider range of benchmarks for semi-supervised learning. For these runs a random class stratified exemplar selection strategy is used.
Method Datasets
EuroSAT DeepWeeds Flowers Pets
PAWS 94.4±0.3 72.6±4.5 98.5±0.7 90.2±1.0
MoLAR-SS 94.2±0.1 87.8±1.5 98.9±0.6 91.6±0.2
$\Leftarrow$ $\Rightarrow$ $\Rightarrow$
(a) PAWS
(b) Random Init.
(c) vMF-SNE
(d) MoLAR-SS
Figure 4: Visualizing the impact of vMF-SNE initialisation. $t$-SNE visualisation of the projection head embeddings at various stages of the training process for DeepWeeds. Blue is the weed present class, and orange is the negative class. The arrows represent the order of training steps.
Table 6: Out of distribution detection ablation. OpenOOD results for MoLAR with and without vMF-SNE initialisation and SKMPS exemplar selection. By default results are reported with these components.
Methods Datasets
CIFAR-10
Near-OOD Far-OOD
MoLAR 91.1 93.0
+vMF-SNE 91.5 94.2
+SKMPS 95.4 97.2
MoLAR-SS 91.4 95.6
+vMF-SNE 93.2 96.6
+SKMPS 93.1 98.0
4 Experiments
4.1 Experimental setup
The experiments closely follow the OpenOOD [46] benchmarks, which use a specific set of near-OOD datasets and far-OOD datasets for each ID dataset. For CIFAR-10, the near-OOD datasets are CIFAR-100 [13] and Tiny ImageNet [15]. The far-OOD images are Places365 [48], DTD [6], MNIST [7] and SVHN [23]. These are the same for CIFAR-100, which instead uses CIFAR-10 as a near-OOD dataset.
Several semi-supervised learning benchmarks are also considered to test the performance of MoLAR-SS. We focus on the well-studied CIFAR-10 (10 classes), CIFAR-100 (100 classes) [13] and Food-101 (101 classes) [3] benchmarks, but also include some contextual results on EuroSAT (10 classes) [9], Flowers-102 (102 classes) [25], Oxford Pets (37 classes) [28] and DeepWeeds (2 classes) [26]. While DeepWeeds has originally 9 classes, we treat it as a binary classification problem to identify weeds versus not weeds, due to the size of the negative class.
When sampling exemplars using SKMPS, USL or using a random class-stratified sample, 4 (CIFAR-10), 4 (CIFAR-100), 4 (EuroSAT), 27 (DeepWeeds - binary), 2 (Flowers-102), 2 (Food-101) and 2 (Oxford Pets) images per class are selected. For accuracy in the semi-supervised learning setting, we report the mean of five runs and the standard deviation, with the same sets of exemplars being used for all experiments unless specified otherwise. For EuroSAT and DeepWeeds, where preset validation splits were not available, we sampled 500 images per class, and 30% of the images for the 9 original classes respectively.
4.2 Results
MoLAR is competitive with previous OOD detection methods.
Tab. 1 shows that MoLAR provides strong performance in comparison to previous methods based on finetuning ResNet backbones, and their application with a frozen DINOv2 backbone. In the latter case a KNN OOD metric [35] is used, which measures distance based on the closest training example in the normalized embedding space. Overall, it appears that CIDER and PALM perform better with a frozen DINOv2 backbone, but they struggle to match the performance of the backbone itself, particularly for near-OOD data. Conversely, MoLAR is able to significantly improve on the near-OOD performance of the backbone in every case, providing particularly strong performance on CIFAR-100.
MoLAR retains strong performance when only using exemplars to detect OOD examples.
Using a KNN metric to identify OOD examples with the whole training dataset provides strong results, but can be computationally inefficient. Using a smaller set of exemplars is much cheaper to compute, and for MoLAR the distance to these exemplars is already computed in the process of making a class prediction. The performance of different OOD detection approaches with a set of exemplars selected using SKMPS is shown in Tab. 2. It is noted that MoLAR retains strong OOD detection performance, and for CIFAR-100 performance actually improves in comparison to Tab. 1. CIDER and PALM do not fare as well in comparison to using the full training dataset for OOD detection, with performance falling in most cases. We also observe strong performance for MoLAR-SS, which performs similarly to MoLAR and provides stronger performance than the fully supervised CIDER and PALM methods.
MoLAR-SS is competitive with previous semi-supervised learning methods.
Tab. 3 shows that across a range of benchmarking datasets MoLAR-SS outperforms PAWS using the same frozen backbone model, and is competitive with other methods in the literature that train models of similar sizes and using the same number (or greater) of labelled images. MoLAR-SS outperforms FreeMatch [42] on CIFAR-10 by 0.8% on average with the smallest DINOv2 model available, and also outperforms Semi-ViT [18] by 5.1% using a much smaller labelled set. SemiReward [18] obtains better results with a similar model size for CIFAR-100, but in contrast to MoLAR this approach fits all model parameters whereas MoLAR-SS uses a frozen backbone. MoLAR-SS outperforms SemiReward (+1.1%) with the ViT-B model, which takes about four hours of GPU time using Nvidia V100 GPUs, compared to the 67 hours of compute time for SemiReward on a more powerful Nvidia A100 GPU [18].
Simple $k$-means exemplar selection (SKMPS) obtains competitive performance in comparison to USL.
Tab. 4 shows that using exemplar selection strategies improves the performance of both PAWS and MoLAR-SS, and that the best performing approach across both methods and all three datasets considered is the SKMPS strategy. While the results between USL [41] and SKMPS are similar for CIFAR-100, significantly better results are obtained on the CIFAR-10 dataset using PAWS, and on the Food-101 dataset using both PAWS and MoLAR-SS. SKMPS also provides more consistent results than USL for the Food-101 dataset, with a smaller variance in performance.
Refer to caption Refer to caption
(a) PAWS
Refer to caption Refer to caption
(b) MoLAR-SS
Figure 5: Visualizing improvements in OOD detection. $t$-SNE visualisation of the projection head embeddings and density plots of minimum cosine distance to an exemplar ($\circ$) for in-distribution CIFAR-10 (blue) and OOD CIFAR-100 (orange).
(a) CIFAR-10
(b) CIFAR-100
(c) Food-101
Figure 6: Visualizing improvements in exemplar selection. $t$-SNE visualisation of the DINOv2 ViT-S/14 backbone embeddings overlayed with the exemplars selected by the USL ($+$) and SKMPS ($\circ$) strategies.
vMF-SNE initialisation can be important for the success of semi-supervised learning with a projection head.
In Tab. 5 the performance of PAWS and MoLAR-SS is explored across a broader range of datasets. In some cases, similar performance between the two approaches is found, while for others there is a drastic difference. This is particularly notable for the DeepWeeds dataset, where the vMF-SNE initialisation is important for obtaining good results. This can be seen in Fig. 4, which shows that in DeepWeeds the randomly initialised projection head clusters the data poorly in the first instance. As a result, training with PAWS results in mediocre performance and a mixing of the two classes (Fig. 4(a)), while training using the vMF-SNE initialised head (Fig. 4(c)) results in better seperation (Fig. 4(d)) and better performance.
SKMPS and vMF-SNE intialisation contribute to OOD detection performance.
Tab. 6 shows that vMF-SNE initialisation and SKMPS contribute to improving OOD performance with both MoLAR and MoLAR-SS.
5 Discussion
MoLAR learns compact representations.
Fig. 5 shows that MoLAR-SS learns a much more compact representation around the labelled exemplars in comparison to PAWS. This can be attributed to MoLAR-SS using an averaging and sharpening approach across views, rather than a consistency based loss, that encourages representations to be more compact while also providing superior semi-supervised learning performance. It is found that (Fig. S8) that the consistency based loss can in fact result in performance degradation over a training run.
SKMPS selects more diversity in comparison to USL.
Fig. 6 shows that SKMPS samples a more diverse set of exemplars in comparison to USL. We explore this further in the supplementary material, showing in Fig. S9 that SKMPS samples more classes than USL on small budgets. For example, for selecting 202 exemplars on the Food-101 dataset, USL never samples all of the classes in 20 runs whereas SKMPS samples all the classes in 25% of cases.
Conclusion.
This paper proposes MoLAR, a unified approach to training supervised and semi-supervised image classifiers that provides state-of-the-art OOD detection results. MoLAR is designed to be used with frozen pretrained foundation models, employing parametric vMF-SNE initialisation and simple $k$-means exemplar selection (SKMPS) to obtain the maximal benefit from the high-quality embeddings that these models make available. This enables significantly shorter training times in comparison to other OOD detection methods that finetune a CNN backbone, and MoLAR can obtain strong OOD detection performance by comparing OOD examples only to a small set of exemplars. This allows for efficient OOD detection, that is no more expensive than making a class prediction. Further, MoLAR obtains competitive semi-supervised learning results in comparison to other approaches designed only to maximize performance.
Acknowledgements
We acknowledge the Traditional Owners of the unceded land on which the research detailed in this paper was undertaken: the Wurundjeri Woi-wurrung and Bunurong peoples. This research was undertaken using the LIEF HPC-GPGPU Facility hosted at the University of Melbourne. This Facility was established with the assistance of LIEF Grant LE170100200. This research was also undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government. Evelyn Mannix was supported by a Australian Government Research Training Program Scholarship to complete this work.
Supplementary Material
Appendix A Parametric vMF-SNE Initialisation
This section describes the parametric Stochastic Neighbour Embedding (SNE) approach [11] used to initialize the projection head of the network in MoLAR. This method involves building two probability distributions that describe the similarity of data points in two different spaces, the origin space $P$ and the target space $Q$. The idea is for similar points $i,j$ in $P$ to have a high probability $p_{ij}$ compared to dissimilar points, and to fit the distribution in the target space $Q$ to the distribution in the origin space $P$ by minimizing the KL divergence
$\text{KL}(Q||P)=\sum_{ij}p_{ij}\log\frac{p_{ij}}{q_{ij}}$ (10)
where $p_{ij}$ and $q_{ij}$ describe the similarity between points in the origin and target space.
These $P,Q$ distributions are built using a probability distribution as a kernel. As MoLAR is defined as a von Mises-Fisher (vMF) mixture model, this is a natural choice. Using the vMF kernel to compute the probabilities $P$ in the latent space defined by our backbone model $f$ provides
$p_{j|i}$ :=$\frac{k_{\kappa_{i}}(\hat{h}_{j};\hat{h}_{i})}{\sum_{k\not=i}k_{\kappa_{i}}(\hat{h}_{k};\hat{h}_{i})},$ (11)
$k_{\kappa_{i}}(\hat{h}_{j};\hat{h}_{i})$ $\propto\text{exp}\left(\hat{h}_{j}\cdot\hat{h}_{i}\,\kappa_{i}\right),$ (12)
where $h_{j}=f(x_{j})$ and we define $p_{i|i}=0$, noting that the vMF normalization constant within $k_{\kappa_{i}}$ cancels out in this expression. We set $\kappa_{i}$ for each point to make the entropy $H_{i}$ equal to the perplexity parameter $\gamma$, $H_{i}=\sum_{k}p_{k|i}\log p_{k|i}=\log\gamma$. This ensures each point has a similar number of neighbours in the distribution $P$, and for each point $\kappa_{i}$ is determined using the bisection method as is standard in SNE approaches. To obtain a symmetric distribution, we use $p_{ij}=\frac{p_{j|i}+p_{i|j}}{2N}$, where $N$ is the number of data points.
To compute the probability distribution $Q$ in the latent space of the projection head, a vMF distribution with fixed concentration $\tau$, as in the main text, is used as the kernel,
$q_{j|i}$ =$\frac{k(\hat{z}_{j};\hat{z}_{i})}{\sum_{k\not=i}k(\hat{z}_{k};\hat{z}_{i})},$ (13)
where $k$ is a MoLAR mixture component (Eq. 2) and we define $q_{i|j}=0$. We also symmetrise this conditional distribution as above to obtain $q_{ij}$.
To minimize the KL divergence in Eq. 10, gradients are taken with respect to the parameters of the projection head $g_{\theta}$. This allows the vMF-SNE initialisation to follow a similar approach as to training the neural networks in MoLAR, utilising mini-batching, stochastic gradient descent and image augmentations [1].
This parametric vMF-SNE initialisation method is a similar to $t$-SNE [38], parametric $t$-SNE [37] and non-parametric vMF-SNE [40], but we apply it in a very different context — as an initialisation approach for neural networks using frozen foundation model embeddings. There are three key hyper-parameters in this process — the batch size, which defines the number of images to sample in each mini-batch, the perplexity parameter $\gamma$, and the concentration of the vMF distributions in $Z_{2}$ which is determined by $\tau$. An ablation study considering the impact of these parameters is presented Sec. C.2.
Algorithm 1 vMF-SNE initialisation
[…]
Randomly initialise MLP head $g_{\theta}$ and set starting value of $\kappa_{i}=1$ for all images $i$ ;
$l=0$;
while $l<N_{E}$ do
     Randomly split $D$ into $B$ mini-batches;
     for $(\textbf{x}_{b},\textbf{y}_{b})\in\{D_{1},...,D_{b},...,D_{B}\}$ do
         Augment $\textbf{X}=\text{Aug}(\textbf{x}_{b})$;
         Calculate $\textbf{H}=f(\textbf{X})$, normalize to obtain $\hat{\textbf{H}}$;
         Calculate $\textbf{Z}=g_{\theta}(\textbf{H})$, normalize to obtain $\hat{\textbf{Z}}$;
         Calculate $k_{\kappa_{i}}(\hat{h}_{j};\hat{h}_{i})=\text{exp}\left(\hat{\textbf{H}}_{j:}\cdot\hat{\textbf{H}}_{i:}\kappa_{i}\right)$ for all images $i,j$;
         Calculate $k(\hat{z}_{j};\hat{z}_{i})=\text{exp}\left(\hat{\textbf{Z}}_{j:}\cdot\hat{\textbf{Z}}_{i:}/\tau\right)$ for all images $i,j$;
         Calculate $p_{j|i}=\frac{k_{\kappa_{i}}(\hat{h}_{j};\hat{h}_{i})}{\sum_{k\not=i}k_{\kappa_{i}}(\hat{h}_{k};\hat{h}_{i})}$;
         Calculate $q_{j|i}=\frac{k(\hat{z}_{j};\hat{z}_{i})}{\sum_{k\not=i}k(\hat{z}_{k};\hat{z}_{i})}$;
         Find optimal values for $\kappa_{i}$ by minimising $\left|\sum_{k}p_{k|i}\log p_{k|i}-\log\gamma\right|$ using bisection method;
         Recalculate $p_{j|i}$ with new values of $\kappa_{i}$;
         Calculate $p_{ij}=\frac{p_{j|i}+p_{i|j}}{2N}$ and $q_{ij}=\frac{q_{j|i}+q_{i|j}}{2N}$;
         Calculate KL divergence $\mathcal{L}=\sum_{ij}p_{ij}\log\frac{p_{ij}}{q_{ij}}$;
         Minimise loss $\mathcal{L}$ by updating $\theta$;
     end for
     $l=l+1$;
end while
Algorithm 2 SKMPS exemplar selection
Input: Array of input images X, $f$ backbone model, labelling budget $k$
Calculate embeddings of all images in the latent space $\textbf{H}=f(\textbf{X})$, and normalize to obtain $\hat{\textbf{H}}$;
Cluster $\hat{\textbf{H}}$ into $k$ clusters using $k$-means to find the cluster centroids C;
Initialise empty list $L=\{\}$;
$i=0$;
while $i<k$ do
     $j=\text{ArgMin}(\text{cosine distance}(\hat{\textbf{H}},\textbf{C}_{i:}))$;
     $L=\{L,j\}$;
     $i=i+1$;
end while
Return: List $L$ of the indices of the images to be labelled
Algorithm 3 MoLAR training algorithm.
[…]
Initialise MLP head $g_{\theta}$ using vMF-SNE intialisation.
$l=0$;
while $l<N_{E}$ do
     Randomly split $D$ into $B$ mini-batches;
     for $(\textbf{x}_{b},\textbf{y}_{b})\in\{D_{1},...,D_{b},...,D_{B}\}$ do
         Select $b_{e}$ exemplars per class from $D_{e}$, to obtain $(\textbf{x}_{e},\textbf{y}_{e})$;
         Augment $\textbf{X}=\text{Aug}(\textbf{x}_{b})$;
         Augment $\textbf{X}_{e}=\text{Aug}(\textbf{x}_{e})$;
         Calculate $\textbf{Z}=g_{\theta}(f(\textbf{X}))$, normalize to obtain $\hat{\textbf{Z}}$;
         Calculate $\textbf{Z}_{e}=g_{\theta}(f(\textbf{X}_{e}))$, normalize to obtain $\hat{\textbf{Z}}_{e}$;
         Calculate $\bm{\phi}=\text{OneHot}(\textbf{y}_{b})(1-\alpha)+\alpha$;
         Calculate $\textbf{P}=\text{Softmax}(\hat{\textbf{Z}}\cdot\hat{\textbf{Z}}_{e}^{\top})\,\bm{\phi}$;
         if labels $\textbf{y}_{b}$ are known then
              Calculate loss $\mathcal{L}=\frac{1}{|D_{b}|}\sum_{i}^{|D_{b}|}H({\textbf{y}_{b}}_{i},\textbf{P}_{i,:})$;
         else
              Split P to obtain the outputs for different views of the same image $\textbf{P}^{*},\textbf{P}^{+}$
              Calculate pseudolabels $\textbf{y}^{*}=\rho(\frac{\textbf{P}^{*}+\textbf{P}^{+}}{2})$;
              Calculate average prediction $\overline{p}=\sum_{i}^{|D|}\rho\left(\frac{\textbf{P}_{i,:}^{*}+\textbf{P}_{i,:}^{+}}{2}\right)$;
              Calculate loss $\mathcal{L}=\frac{1}{2|D_{b}|}\sum_{i}^{|D_{b}|}\left(H({\textbf{y}_{i,:}^{*}},\textbf{P}_{i,:}^{+})+H({\textbf{y}_{i,:}^{*}},\textbf{P}_{i,:}^{*})\right)-H(\overline{p})$;
         end if
         Minimise loss $\mathcal{L}$ by updating $\theta$;
     end for
     $l=l+1$;
end while
Appendix B Additional results
Figure S7: Semi-supervised learning ablation study. Comparison of the vMF-SNE initialisation and the averaging and sharpening (MixMatch) loss components of MoLAR-SS versus the original PAWS (Consistency) approach.
Semi-supervised learning ablation studies.
Fig. S7 shows that across the CIFAR-10, CIFAR-100, Food-101 and DeepWeeds datasets, using MoLAR-SS with vMF-SNE initialisation leads to the best overall results. For CIFAR-10 and CIFAR-100, the averaging and sharpening (MixMatch) loss significantly improves overall accuracy and the gain from vMF-SNE initialisation is less obvious. However, for DeepWeeds skipping vMF-SNE initialisation leads to poor results. In the case of the Food-101 dataset, the combination of both vMF-SNE initialisation and the MixMatch loss leads to be best average accuracy across all training runs.
Figure S8: Training progress curves. Training curves comparing the vMF-SNE and random initialisation of the projection head and consistency versus the MoLAR-SS (Mixmatch) loss.
Consistency based loss can degrade the performance of PAWS during a training run.
In Fig. S8 the training curves for two of the datasets from Fig. S7 are shown. These training curves demonstrate that the consistency loss as used in PAWS can result in performance degrading for CIFAR-10 after around 10 epochs. When using the averaging and sharpening (MixMatch) loss as done in MoLAR-SS, the validation accuracy remains stable once the models have converged. This is likely caused when different views on the class boundaries have highly confident, but different, sharpened class probabilities. For PAWS this would result in different views of the same image being given different labels during training, resulting in instabilities at the class boundary. However, in MoLAR-SS this would result in one class or the other being chosen, or a more uncertain result, providing a more consistent supervision signal.
It is also noted that the performance improvement from vMF-SNE initialisation is not due to a head-start in training. While it does provide a better starting point for learning, in both cases models are trained until they completely converge, and vMF-SNE initialised projection heads converge to higher performing maxima, particularly for the DeepWeeds dataset where training using a randomly initialised projection head leads to poor performance.
(a) Number of classes. Average number of classes sampled over 20 runs. Error bars show maximum and minimum number of classes sampled.
(b) Proportion of classes. Proportion of sets sampling all classes over 20 runs.
Figure S9: Exemplar selection diversity. Class diversity captured by different exemplar selection strategies.
SKMPS selects a more diverse set of images in comparsion to USL.
In Fig. S9 we compare diversity of classes sampled for the USL and SKMPS strategies for progressively larger budgets for the CIFAR-10, CIFAR-100 and Food-101 datasets. In all three cases, the SKMPS strategy samples a greater diversity of images, as measured by the number of classes sampled from each dataset for a given budget. When not all classes are sampled, this reduces the maximum attainable performance for a model as classes that are not represented by atleast one labelled image cannot be learned in most semi-supervised learning frameworks. The USL approach considers density to be an indicator of exemplar informativeness, and preferentially chooses exemplars from denser regions of the embedding space of the foundation model [41]. In contrast, SKMPS only considers the density of the points through the $k$-means clustering process, which results in a more diverse sample in addition to providing a more computationally efficient approach.
Table S7: Semi-supervised OOD detection with exemplars: Near OOD. Detection performance on near OOD datasets with the DINOv2 ViT-S/14 frozen backbone. The $\uparrow$ means larger values are better and the $\downarrow$ means smaller values are better.
IDD Method Near OOD Datasets Average
CIFAR-10 CIFAR-100 Tiny ImageNet
[…]
CIFAR-100 PAWS 90.11 51.13 90.99 46.54 90.55 48.84
MoLAR-SS 95.79 26.05 94.60 32.30 95.19 29.18
backbone 85.51 60.16 91.13 30.6 88.32 45.38
CIFAR-10 PAWS 82.10 73.90 83.19 69.49 82.64 71.70
MoLAR-SS 91.61 43.36 91.12 42.78 91.37 43.07
+vMF-SNE 92.84 40.43 93.50 34.92 93.17 37.68
+SKMPS 93.34 35.94 92.83 33.30 93.09 34.62
backbone 90.66 40.13 97.16 11.12 93.91 25.62
Table S8: Semi-supervised OOD detection with exemplars: Far OOD. Detection performance on far OOD datasets with the DINOv2 ViT-S/14 frozen backbone. The $\uparrow$ means larger values are better and the $\downarrow$ means smaller values are better.
IDD Method Far OOD Datasets Average
DTD MNIST SVHN Places365
[…]
CIFAR-100 PAWS 93.59 35.71 97.30 13.95 94.33 32.66 92.78 39.65 94.50 30.49
MoLAR-SS 96.55 22.29 98.96 2.27 94.35 33.39 95.22 28.58 96.27 21.63
backbone 98.92 4.8 87.98 86.15 72.76 85.57 97.03 12.18 89.17 47.18
CIFAR-10 PAWS 91.36 48.71 96.44 15.37 84.99 81.74 89.27 54.24 90.52 50.01
MoLAR-SS 95.44 27.84 98.67 5.35 93.81 39.49 94.48 27.4 95.6 25.02
+vMF-SNE 96.05 24.95 98.68 1.91 95.50 28.77 96.06 22.20 96.57 19.46
+SKMPS 97.45 15.62 99.78 0.00 98.04 10.78 96.70 17.13 97.99 10.88
backbone 99.96 0.16 99.4 1.91 80.57 71.18 98.8 4.17 94.68 19.36
The MoLAR-SS and MoLAR projection heads provides more reliable OOD detection than the backbone embedding.
In Tab. S7 and Tab. S8 we present the full set of results for each dataset in the OpenOOD benchmark, for the semi-supervised learning approaches that were considered. In these tables, only the exemplars that were selected (either using a random class stratified selection or SKMPS) were used to determine if an image was OOD. Results are also provided for the backbone embedding. In general we find that the backbone embedding performs better on the datasets that are included in the DINOv2 training set, such as ImageNet, DTD and Places365 (which is included in the ADE20K dataset [49]). However, the backbone performs very poorly on OOD detection for datasets it hasn’t seen before such as CIFAR-10, CIFAR-100, MNIST and SVHN, while the MoLAR-SS projection head can still perform well. Overall, the backbone performs slightly better on the near OOD CIFAR-10 benchmark (+0.82), while the MoLAR-SS embeddings perform better on the far OOD CIFAR-10 (+3.31) and both CIFAR-100 (+6.87/+7.10) benchmarks. Similar trends are observed for the supervised cases in Tab. S9, Tab. S10, Tab. S11 and Tab. S12.
Table S9: Supervised OOD detection: Near OOD. Detection performance on near OOD datasets with the DINOv2 ViT-S/14 frozen backbone. The $\uparrow$ means larger values are better and the $\downarrow$ means smaller values are better.
IDD Method Near OOD Datasets Average
CIFAR-10 CIFAR-100 Tiny ImageNet
[…]
CIFAR-100 CIDER 76.7 79.32 76.35 76.72 76.52 78.02
PALM 87.32 55.58 85.44 55.49 86.38 55.54
MoLAR 95.08 30.15 95.93 22.7 95.5 26.42
backbone 87.97 56.05 91.82 29.63 89.9 42.84
CIFAR-10 CIDER 93.88 33.38 90.96 32.46 92.42 32.92
PALM 95.13 24.46 93.47 25.02 94.3 24.74
MoLAR 95.17 23.29 98.39 6.68 96.78 14.99
backbone 94.08 29.27 97.58 10.3 95.83 19.78
Table S10: Supervised OOD detection: Far OOD. Detection performance on far OOD datasets with the DINOv2 ViT-S/14 frozen backbone. The $\uparrow$ means larger values are better and the $\downarrow$ means smaller values are better.
IDD Method Far OOD Datasets Average
DTD MNIST SVHN Places365
[…]
CIFAR-100 CIDER 89.05 56.77 99.92 0.0 97.4 15.23 78.28 77.36 91.16 37.34
PALM 94.74 29.93 99.8 0.0 97.19 16.12 90.88 43.46 95.65 22.38
MoLAR 97.59 14.34 99.3 2.1 94.7 28.95 95.83 23.04 96.86 17.11
backbone 97.89 8.48 92.25 52.51 82.32 82.49 96.19 15.95 92.16 39.86
CIFAR-10 CIDER 97.25 14.4 99.58 0.04 99.3 1.07 94.07 22.99 97.55 9.62
PALM 98.34 7.57 99.98 0.0 99.73 0.45 96.5 14.04 98.64 5.52
MoLAR 98.71 5.39 99.99 0.0 95.72 16.98 98.34 6.77 98.19 7.29
backbone 99.96 0.18 99.88 0.01 93.65 39.63 98.88 4.45 98.09 11.07
Table S11: Supervised OOD detection with exemplars: Near OOD. Detection performance on near OOD datasets with the DINOv2 ViT-S/14 frozen backbone. The $\uparrow$ means larger values are better and the $\downarrow$ means smaller values are better.
IDD Method Near OOD Datasets Average
CIFAR-10 CIFAR-100 Tiny ImageNet
[…]
CIFAR-100 CIDER 79.8 67.93 77.32 71.92 78.56 69.92
PALM 86.58 57.54 83.73 59.41 85.16 58.48
MoLAR 95.58 29.37 96.27 23.08 95.93 26.22
backbone 85.51 60.16 91.13 30.6 88.32 45.38
CIFAR-10 CIDER 92.51 44.27 88.87 46.74 90.69 45.5
PALM 92.79 42.25 90.53 44.52 91.66 43.38
MoLAR 89.82 47.86 92.29 37.45 91.05 42.66
+vMF-SNE 89.26 51.06 93.73 31.61 91.5 41.34
+SKMPS 93.83 33.79 96.97 15.64 95.4 24.72
backbone 90.66 40.13 97.16 11.12 93.91 25.62
Table S12: Supervised OOD detection with exemplars: Far OOD. Detection performance on far OOD datasets with the DINOv2 ViT-S/14 frozen backbone. The $\uparrow$ means larger values are better and the $\downarrow$ means smaller values are better.
IDD Method Far OOD Datasets Average
DTD MNIST SVHN Places365
[…]
CIFAR-100 CIDER 87.13 60.92 99.85 0.01 95.91 22.88 81.65 64.57 91.14 37.1
PALM 93.08 39.43 99.71 0.09 95.82 23.47 89.41 48.27 94.51 27.82
MoLAR 97.14 18.94 99.37 0.62 95.34 29.4 96.22 23.64 97.02 18.15
backbone 98.92 4.8 87.98 86.15 72.76 85.57 97.03 12.18 89.17 47.18
CIFAR-10 CIDER 95.77 27.23 99.98 0.0 99.25 1.23 92.15 36.04 96.79 16.13
PALM 95.84 28.37 99.78 0.0 99.29 2.56 93.55 34.03 97.11 16.24
MoLAR 94.21 34.29 96.83 15.45 88.78 48.0 91.97 39.08 92.95 34.2
+vMF-SNE 93.94 35.55 99.14 0.88 89.93 45.45 93.58 32.69 94.15 28.64
+SKMPS 96.98 17.3 99.79 0.0 94.87 26.06 97.04 15.18 97.17 14.64
backbone 99.96 0.16 99.4 1.91 80.57 71.18 98.8 4.17 94.68 19.36
Table S13: Transfer learning with vMF-SNE. Transferability of vMF-SNE initialisation between datasets.
Method Components Dataset
Head init. DeepWeeds
MoLAR-SS Random init. 76.0±9.3
vMF-SNE init. on C100 79.9±8.1
vMF-SNE init. on C10 81.5±7.1
vMF-SNE init. on Food 84.2±6.0
vMF-SNE init. on DeepWeeds 87.8±1.5
Projection heads initialised using the vMF-SNE approach perform better than random initialisation, even if they are initialised using a different dataset.
In (Tab. S13) we fit MoLAR-SS models on the DeepWeeds dataset randomly initialising the projection head, and then use projection heads initialised on various other datasets using the vMF-SNE approach. We found that initialising the projection head with vMF-SNE is better than random initialisation in every case, but best performance was obtained when initialisation was undertaken on DeepWeeds. Potentially, the vMF-SNE approach could be used as a method to initialise a projection head for a foundational model on a large dataset, and this could then be used to obtain better performance than random initialisation on smaller datasets.
Table S14: RoPAWS. Compatability of a MixMatch inspired loss and vMF-SNE initialisation with RoPAWS.
Method Components Dataset
Head init. Loss type Label selection C10 DeepWeeds
RoPAWS [22] Random init. RoPAWS-Consistency Random stratified 93.3±1.5 81.2±6.8
RoPAWS-VM vMF-SNE RoPAWS-Mixmatch Random stratified 95.7±0.1 86.4±0.6
vMF-SNE initialisation and a Mixmatch inspired loss also provide performance improvements when used with RoPAWS.
It is shown in Tab. S14 that the improvements introduced in MoLAR-SS for learning with a frozen backbone model also improves the performance obtained when using RoPAWS [22] on in-distribution data.
Table S15: Adding terms to increase class separation. Adding terms from the PALM [20] loss that improve OOD detection performance to PAWS [1] and MoLAR-SS.
Method Components Dataset
PALM [20] $\mathcal{L}_{\text{proto-contra}}$ C10
PAWS Y 92.1±0.1
PAWS N 92.9±0.2
MoLAR-SS Y 94.2±0.1
MoLAR-SS N 95.9±0.1
Naively incoporating elements of the PALM loss into PAWS and MoLAR-SS harms SSL performance.
It is shown in Tab. S15 that incorporating the loss term from PALM that encourages exemplars within a class to be compact, and far from exemplars of other classes, given by
$\mathcal{L}_{\text{proto-contra}}=-\frac{1}{CK}\sum_{c=1}^{C}\sum_{k=1}^{K}\log\frac{\sum_{k^{\prime}=1}^{K}\mathbb{1}(k^{\prime}\not=k)\exp{(\hat{z}_{c}^{k^{\prime}}\cdot\hat{z}_{c}^{k\top}/\tau)}}{\sum_{c^{\prime}=1}^{C}\sum_{k^{\prime\prime}=1}^{K}\mathbb{1}(k^{\prime\prime}\not=k,c^{\prime}\not=c)\exp{(\hat{z}_{c^{\prime}}^{k^{\prime\prime}}\cdot\hat{z}_{c^{\prime}}^{k\top}/\tau)}}$ (14)
results in poorer performance in comparison to not adding this loss term. Here $\mathbb{1}(\cdot)$ is an indicator function that avoids contrasting between the same exemplar, and $\hat{z}_{c}^{k}$ refers to the $k$th exemplar of class $c$, with $K$ exemplars belong to one of the $C$ classes.
Appendix C Additional results for vMF-SNE initialisation
C.1 Performance
Table S16: vMF-SNE intialisation performance. KNN accuracy of the model backbone versus the vMF-SNE initialised projection head.
Embedding Dataset
C10 C100 EuroSAT DeepWeeds Flowers Food Pets
DINOv2 ViT-S/14 backbone 96.2 82.3 90.0 88.7 99.3 80.9 91.5
vMF-SNE proj. head 96.3±0.1 79.0±0.1 91.9±0.1 91.3±0.1 98.2±0.5 81.0±0.1 93.0±0.1
In a classification context, the performance of the backbone model can be measured non-parametrically using a weighted k-Nearest Neighbor (kNN) Classifier [44]. This provides for a quantitative approach to measure the performance of the projection head initialised using vMF-SNE, as reproducing the local structure of the backbone model should provide a similar kNN accuracy. We find that for five of the seven datasets considered, we can meet or beat the kNN accuracy of the backbone model through vMF-SNE initialisation (Tab. S16).
C.2 Hyperparameter ablation study
Table S17: vMF-SNE initialisation ablation study. The results for the parameters used in the other results in this paper are reported as the mean KNN accuracy and standard deviation of five runs, while the other cases report a single run.
Perplexity ($\gamma$) C10 C100 Food
5 96.1 74.0 81.7
30 96.3±0.1 79.0±0.1 81.0±0.1
50 96.3 79.5 80.9
100 96.3 80.1 80.7
vMF concentration ($\tau$) C10 C100 Food
0.01 96.4 79.4 78.8
0.10 96.3±0.1 79.0±0.1 81.0±0.1
1.00 91.7 54.9 68.5
$Z_{2}$ dimension C10 C100 Food
128 96.3 79.1 81.0
256 96.3 79.0 81.0
384 96.4 79.1 81.0
512 96.3±0.1 79.0±0.1 81.0±0.1
1024 96.3 79.0 81.0
Batch size C10 C100 Food
128 96.2 79.5 80.4
256 96.3 79.1 80.7
512 96.3±0.1 79.0±0.1 81.0±0.1
1024 96.3 79.2 81.3
Kernel C10 C100 Food
vMF 96.3±0.1
Gaussian 91.4±0.6
We undertook an ablation study to determine the sensitivity of vMF-SNE initialisation performance to the hyper-parameters within the algorithm. It was found that the method was highly insensitive to most parameters for the CIFAR-10 dataset, with only a significantly lower perplexity $\gamma$ or an order of magnitude higher concentration parameter $\tau$ resulting in changes to the kNN validation accuracy outside a single standard deviation (Tab. S17). The optimal parameter set appeared to be different between CIFAR-100 and the Food-101 datasets, with better performance for CIFAR-100 with a larger perplexity and smaller concentration and batch size, but better performance for Food-101 with a smaller perplexity and larger batch size. We also explored using a Gaussian kernel with CIFAR-10, and found that this resulted in much poorer performance.
Appendix D Further implementation details
D.1 MoLAR, PAWS, CIDER and PALM
We adapted the PAWS and RoPAWS PyTorch [29] implementation available at this GitHub repository to our own codebase using PyTorch-Lightning [8] and Hydra [45], to enable fair comparisons between PAWS and MoLAR-SS. We further adapted the official CIDER and PALM GitHub implementations to our codebase, and used them as the basis for the CIDER and PALM results presented here with a DINOv2 backbone to compare to MoLAR.
All of the MoLAR, MoLAR-SS and PAWS models that are fitted in this paper were trained for 50 epochs, and the vMF-SNE initialisation runs were trained for 20 epochs. This meant that even for the Food-101 dataset with 75,000 images, training a MoLAR or MoLAR-SS model took less than two hours with two Nvidia V100 32GB graphics cards for the ViT-S/14 and ViT-B/14 backbone, and around six hours for the ViT-L/14 backbone. For fair comparisons to the CIDER and PALM approaches, we trained these for 70 epochs to account for the additional vMF-SNE initialisation steps in MoLAR.
D.2 vMF-SNE initialisation
We re-used the same augmentations, optimizer, scheduler and other details as for MoLAR and MoLAR-SS. The key difference is that the vMF-SNE initialisation approach only requires the unlabelled dataloader, compared to MoLAR-SS which requires both a exemplar and unlabelled dataloader. This meant vMF-SNE initialisation was slightly faster than training. We note than when using image augmentations for this approach, we treat each augmented view independently.
D.3 $k$-means clustering and USL
We used the GPU-optimised $k$-means method available within cuML [32]. We save the normalised global class token from the DINOv2 model applied to each image of the dataset using the validation transformations, and run the $k$-means clustering method over this matrix. This method is extremely fast and can easily process millions of rows. We use the best clustering result from ten runs, and choose the number of clusters according to the number of images we wish to label. To select the exemplar to label from each cluster, we choose the image with the largest cosine similarity to the cluster centroid.
To reproduce the USL [41] approach, we used the code published by the authors on GitHub. There are quite a few hyperparameters that are used within the USL method, so we selected the set that was used for ImageNet with CLIP [30], another foundation model based on the ViT architecture, for all of our experiments.